### v4

Abstract
This paper introduces a computationally efficient method that converges globally to B-stationary points of mathematical programs with equilibrium constraints (MPECs). B-stationarity is necessary for optimality and means that no feasible first-order direction can improve the objective. It can be certified by solving a linear program with equilibrium constraints (LPEC) constructed at a given feasible point. The proposed method solves a finite sequence of LPECs, which either certify B-stationarity or provide an active-set estimate for the complementarity constraints, along with branch nonlinear programs (BNLPs) obtained by fixing the complementarity active set in the MPEC. In particular, the method proceeds in two phases: the first identifies a feasible BNLP or a stationary point of a constraint infeasibility minimization problem, and the second solves a sequence of BNLPs until a B-stationary point of the MPEC is found. We prove that under the MPEC-MFCQ, the method requires solving only a finite number of BNLPs and LPECs for convergence. Moreover, we show that, unless the current iterate is B-stationary, the combinatorial LPECs need not be solved to optimality. For convergence, it suffices to compute a nonzero feasible point, which in practice often requires solving a single linear program, yielding significant computational savings. Numerical experiments show that the proposed method is more robust and faster than relaxation-based methods and mixed-integer NLP reformulations (which, in contrast to the proposed approach, do not provide a certificate of B-stationarity), even on medium- to large-scale instances.
1 Introduction
We study mathematical programs with equilibrium constraints (MPECs) of the following form:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $f(x)$ (1a)
s.t. $c(x)\geq 0,$ (1b)
$0\leq x_{1}\perp x_{2}\geq 0,$ (1c)
with the partition of variables $x=(x_{0},x_{1},x_{2})\in{\mathbb{R}}^{n}$, with $x_{0}\in{\mathbb{R}}^{p}$, $x_{1},x_{2}\in{\mathbb{R}}^{m}$. The functions $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $c:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{c}}$ are assumed to be at least twice continuously differentiable. The notation $0\leq x_{1}\perp x_{2}\geq 0$ means that $x_{1,i},x_{2,i}\geq 0$ and $x_{1,i}x_{2,i}=0$ for all $i$. There is no loss of generality in using (1c), since the more general form $0\leq g(x)\perp h(x)\geq 0$, with $g,h:{\mathbb{R}}^{p}\to{\mathbb{R}}^{m}$ twice continuously differentiable, can be transformed into (1c) by introducing slack variables $x_{1}=g(x)$ and $x_{2}=h(x)$. For lighter notation we omit equality constraints in (1). If given by twice continuously differentiable functions, all results extend directly to this case.
Equilibrium constraints are more general than complementarity constraints (1c), but coincide under suitable conditions Luo1996 . We focus on complementarity constraints, yet follow the literature in using the acronym MPEC.
Solving MPECs reliably and fast is of great practical interest, as they arise in numerous applications including process engineering Baumrucker2008 , robotics Wensing2023 , optimal control of hybrid dynamical systems Nurkanovic2023f ; Nurkanovic2022b , bi-level optimization Kim2020 , nonsmooth optimization Hegerhorst2020 , inverse optimization Albrecht2017 ; Hu2012 , to name a few examples. A related class of problems of interest, equivalent to MPEC, are mathematical programs with vanishing constraints Achtziger2008 .
The complementarity constraints (1c) complicate the theoretical and computational aspects of the optimization problem (1). These constraints can be replaced by a set of inequality constraints, leading to a standard Nonlinear Program (NLP):
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $f(x)$ (2a)
s.t. $c(x)\geq 0,$ (2b)
$x_{1,i}\geq 0,\ x_{2,i}\geq 0,\ x_{1,i}x_{2,i}\leq 0,\ \forall i\in\{1,\ldots,m\}.$ (2c)
However, this NLP is numerically and theoretically degenerate. Standard constraint qualifications such as the Mangasarian-Fromovitz constraint qualification (MFCQ) fail at all feasible points Jane2005 ; Scheel2000 , making standard NLP methods inefficient Kim2020 ; Nurkanovic2024b due to unbounded Lagrange multiplier sets Gauvin1977 . Moreover, standard KKT conditions may not apply to (2), complicating stationary point definition and certification Scheel2000 . Various reformulations of (1c) exist, including smooth and nonsmooth C-functions Facchinei2003 ; Leyffer2006a , but none resolve the inherent degeneracy of NLP reformulations of MPECs.
A necessary optimality condition is the absence of feasible first-order descent directions. In the MPEC literature, such points are called B-stationary points. For regular NLPs, B-stationary points simply satisfy KKT conditions, but directly solving KKT conditions for (2) may be difficult or impossible due to degeneracy. Weaker stationarity notions exist, but most allow first-order descent directions, making them too weak as practical stopping criteria Leyffer2007 .
To cope with the degeneracy of NLP reformulations of MPECs such as (2), the optimization community has also developed MPEC-tailored algorithms in the attempt to characterize and compute local minimizers of (1), cf. Hoheisel2013 ; Kanzow2015 ; Kim2020 ; Luo1996 ; Nurkanovic2024b for surveys. We briefly survey some of these methods. They can be divided into regularization and active set/combinatorial methods.
Regularization methods replace the difficult constraints  (2c) with a smoothed or relaxed version of them. For example, in Scholtes’ global relaxation Scholtes2001 $x_{1,i}x_{2,i}\leq 0$ is replaced by $x_{1,i}x_{2,i}\leq\tau$ for all $i$ and $\tau>0$. This yields a regular NLP that usually satisfies MFCQ and can be solved by standard NLP algorithms. In practice, a sequence of NLPs with decreasing $\tau$ is solved to approximate the MPEC solution. Many refinements of this idea exist DeMiguel2005 ; Kadrani2009 ; Kanzow2013 ; Lin2003 ; Scholtes2001 ; Steffensen2010 . Penalty methods Anitescu2005 ; Anitescu2005a ; Ralph2004 handle $x_{1,i}x_{2,i}\leq 0$ by adding a penalty term to the objective, e.g. $\rho\sum_{i=1}^{m}x_{1,i}x_{2,i}$. A sequence of NLPs with increasing $\rho>0$ is then solved to approximate a solution. Under suitable conditions, even finite $\rho$ may yield an MPEC stationary point via a standard NLP Anitescu2005 ; Ralph2004 . However, unless restrictive assumptions hold, all these methods may converge to points that are not B-stationarity Hoheisel2013 ; Kanzow2015 .
Active set methods take a different approach to obtaining a sequence of regular NLPs, whose solutions eventually solve the MPEC. The main idea is to fix the active set complementarity constraints, i.e., replace (1c) by $x_{1,i}=0,x_{2,i}\geq 0,\forall i\in{\mathcal{I}}_{1}$ and $x_{1,j}\geq 0,x_{2,j}=0,\forall j\in{\mathcal{I}}_{2}$, where ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ are a partition of $\{1,\ldots,m\}$. The resulting NLPs – called branch NLPs (BNLPs) – are now regular problems that can be solved with classical NLP methods. Given the correct partition, the solution of such an NLP also solves the MPEC (1). However, finding such a partition and certifying optimality has combinatorial complexity Scheel2000 . Most active set methods assume that MPEC-LICQ holds (cf. Def. 2.2) and search for partitions by looking at the Lagrange multipliers of the active complementarity constraints to either find a new active set or to certify B-stationarity. Such methods are reported in Fukushima2002 ; Giallombardo2008 ; Izmailov2008 ; Jiang1999 ; Lin2006 ; Liu2006 ; Luo1996 ; Scholtes1999 . However, if MPEC-LICQ does not hold, only weaker concepts of stationarity – possibly allowing descent directions – can be used as stopping criteria.
To avoid the restrictive assumption of MPEC-LICQ, B-stationarity can be verified by solving a linear program with equilibrium constraints (LPEC) Luo1996 , obtained by linearizing the functions $f$ and $c$ in (1), but retaining the complementarity constraints. If the current point is not B-stationary then the LPEC provides a descent direction and a corresponding active set guess. Such an approach is taken in Guo2022 ; Kirches2022 ; Leyffer2007 . Leyffer and Munson Leyffer2007 propose a sequential LPEC QP method. Since solving an LPEC can be expensive, Kirches et al. Kirches2022 study simpler LPECs with only complementarity and box constraints, and suggest handling of other constraints via an augmented Lagrangian. Guo and Deng Guo2022 analyze this approach and prove convergence to M-stationary points (Def. 2.3). However, M-stationarity may still admit descent directions, cf. (Leyffer2007, , Sec 2.1). A main drawback of these methods is their implementation complexity, since all components of a robust solver must be provided.
Finally, we mention the class of hybrid methods Kazi2024 ; Lin2006 , which combine active set and regularization-based methods. The method presented here, which we call MPEC optimizer (MPECopt), can be classified as such a method and consists of two phases. Phase I finds a feasible BNLP by solving relaxed MPECs for solution approximations, then LPECs to identify valid active sets. Phase II starts from a feasible BNLP and solves LPECs at each BNLP’s stationary point until reaching B-stationarity. LPEC solutions either certify B-stationarity or provide active set guesses for BNLPs with better objective values. Notably, the LPECs steps never update iterates, but only provide better active sets or certify B-stationarity.
We highlight key differences from other hybrid approaches. Unlike Lin and Fukushima Lin2006 , our method solves LPECs for active set guesses rather than using multipliers, eliminating MPEC-LICQ requirements and allowing convergence to B-stationary points that are not S-stationary. Unlike Kazi et al. Kazi2024 , we never apply LPEC steps directly, ensuring all Phase II iterates remain feasible.
It is important to note that LPECs are nonconvex combinatorial problems whose efficient solution is crucial for our method’s effectiveness. LPECs with only bound and complementarity constraints have linear complexity (Kirches2022, , Propositions 2.3 and 2.4). As special cases of MPECs, LPECs can be solved with regularization and penalty methods. Tailored LPEC methods include simplex extensions Fang2010 , mixed-integer linear programming (MILP) approaches with tailored cuts Hu2012 ; JaraMoroni2020 , and specialized penalty methods Hall2024 ; JaraMoroni2018 ; see Judice2012 for a survey. However, regularization and penalty methods may converge to points with trivial descent directions and fail to verify B-stationarity, making combinatorial methods preferable. Importantly, we show that, in our proposed method, LPECs need only be solved to optimality when certifying a B-stationary point at the last iteration of Phase II. Otherwise, in Phase I, it is sufficient to find a feasible point of the LPEC, and in Phase II, a feasible descent direction. Our numerical experiments show that this is often equivalent to solving a single LP.
Contributions: This paper presents a new method, called MPEC optimizer (MPECopt), that computes B-stationary points by solving a finite number of LPECs and NLPs in two phases. We prove that, by finding a feasible point of an LPEC constructed at a point sufficiently close to the feasible set of the MPEC, the active set obtained from the LPEC corresponds to a BNLP that is guaranteed to be feasible. For our implementation, this result guarantees that a Phase I, which combines well-established regularization methods and an LPEC solver, finds feasible points of a BNLP (in fact, already stationary points of the MPEC). Phase II either certifies B-stationarity of this point or iterates over a sequence of feasible BNLPs while strictly decreasing the objective until a B-stationary point is found. We provide a detailed convergence analysis and prove that, under the MPEC-MFCQ, Phase II is guaranteed to find a B-stationary point in a finite number of LPEC and BNLP solves.
Furthermore, we discuss in detail two different LPEC formulations, their relation, and theoretical properties. We show that, unless we need to certify B-stationarity, it is sufficient to find a feasible (improving) point of the LPEC to ensure progress, which often reduces the computational effort significantly. In practice, solving a single linear program is often sufficient. We also discuss in detail why LPECs can be solved efficiently via their MILP reformulation.
We compare the proposed method with several relaxation-based methods and a mixed-integer nonlinear programming (MINLP) reformulation on two benchmarks: the MacMPEC set and a synthetic large-scale nonlinear MPEC benchmark. Higher success rates and generally faster computation times are reported. Interestingly, solving LPECs turns out not to be a computational bottleneck in practice, even for large problem instances. We highlight that most other methods with an available implementation – including those we compare to – do not, in general, provide a certificate of B-stationarity for generic MPECsand they may converge to a point with a first-order descent direction. Additionally, an open-source implementation of the method is provided at github.com/nosnoc/mpecopt.
Notation: Subscripts denote vector/matrix components ($x_{i}$), superscripts denote iterates ($x^{k}$). Functions evaluated at particular iterates are denoted as $f^{k}:=f(x^{k})$. The concatenation of two vectors $a\in R^{n}$ and $b\in{\mathbb{R}}^{m}$ is shortly written as $(a,b)=\begin{bmatrix}a^{\top}&b^{\top}\end{bmatrix}^{\top}$. The same notation is adapted accordingly to the concatenation of several vectors. For scalars $a,b\in{\mathbb{R}}$, $\max(a,b)$ returns the larger value. For vectors $a,b\in{\mathbb{R}}^{n}$: $\max(a)$ returns the largest component, $c=\max(a,b)$ gives $c_{i}=\max(a_{i},b_{i})$. The $\min$ function is defined in all cases accordingly. Let $x\in{\mathbb{R}}^{n}$, then $\mathrm{diag}(x)\in{\mathbb{R}}^{n\times n}$ returns a diagonal matrix with $x_{i}$ as its diagonal elements.
Outline: Section 2 recalls some preliminaries. Section 3 provides a detailed statement of MPECopt, discusses how to formulate and solve LPECs. Section 4 gives a detailed convergence analysis of the new method. Section 5 provides all implementation details and numerical results on two benchmark test sets. Section 6 provides concluding remarks.
2 Preliminaries
We start by defining some useful index sets for the complementarity constraints of the MPEC (1). Denote the feasible set of the MPEC (1) by $\Omega=\{x\in{\mathbb{R}}^{n}\mid c(x)\geq 0,0\leq x_{1}\perp x_{2}\geq 0\}$. The following index sets, which depend on a feasible point $x$ define a partition of $\{1,\ldots,m\}$:
${\mathcal{I}}_{0+}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}>0\},$
${\mathcal{I}}_{+0}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}>0,x_{2,i}=0\},$
${\mathcal{I}}_{00}(x)$ $=\{i\in\{1,\ldots,m\}\mid x_{1,i}=0,x_{2,i}=0\}.$
Most of the computational and theoretical difficulties arise if the so-called biactive set ${\mathcal{I}}_{00}(x)$ is nonempty. The active set of the standard inequality constraints (1b) is defined as ${\mathcal{A}}(x)=\{i\in\{1,\ldots,n_{c}\}\mid c_{i}(x)=0\}$}.
Next, we state first-order necessary optimality conditions. For this we need the notion of the tangent cone at $x\in\Omega$ to the set $\Omega$, which is defined as $\mathcal{T}_{\Omega}(x)=\{d\in{\mathbb{R}}^{n}\mid\exists\{x^{k}\}\subset\Omega,\{t^{k}\}\subset{\mathbb{R}}_{\geq 0}:\lim\limits_{k\to\infty}t^{k}=0,\lim\limits_{k\to\infty}x^{k}=x,\lim\limits_{k\to\infty}\frac{x^{k}-x}{t^{k}}=d\}$.
Theorem 2.1
((Luo1996, , Corollary 3.3.1)) Let $x^{*}\in\Omega$ be a local minimizer of (1), then it holds that
$\nabla f(x^{*})^{\top}d\geq 0\;\textrm{ for all }d\in\mathcal{T}_{\Omega}(x^{*}),$ (3)
or equivalently, $d=0$ is a global optimizer of the following optimization problem:
$\min_{d\in{\mathbb{R}}^{n}}\quad$ $\nabla f(x^{*})^{\top}d\quad\textnormal{s.t.}\quad d\in\mathcal{T}_{\Omega}(x^{*}).$ (4)
If a point $x^{*}$ satisfies the condition above, it is said that geometric Bouligand stationarity holds, or for short, $x^{*}$ is geometric B-stationary Flegel2005a ; Luo1996 ; Scheel2000 .
Note that Theorem 2.1 is purely geometric and thus difficult to verify. In standard optimization theory, to obtain an algebraic characterization of a stationary point, the tangent cone ${\mathcal{T}}_{\Omega}(x)$ is replaced by the linearized feasible cone (Nocedal2006, , Def. 12.3). However, the latter is always a convex polyhedral cone, and if $\mathcal{I}_{00}(x)$ is nonempty the cone $\mathcal{T}_{\Omega}(x)$ is always nonconvex Luo1996 , making this transition impossible. A more suitable notion is the so-called MPEC-linearized feasible cone Flegel2005a ; Pang1999 ; Scheel2000 , which for a feasible point $x\in\Omega$ reads as
$\begin{split}\mathcal{F}_{\Omega}(x)=\{d\in{\mathbb{R}}^{n}\mid\ &c_{i}(x)+\nabla c_{i}(x)^{\top}d\geq 0,\ \forall i\in{\mathcal{A}}(x),\\ &x_{1,i}+d_{1,i}=0,\ \forall i\in\mathcal{I}_{0+}(x),\\ &x_{2,i}+d_{2,i}=0,\ \forall i\in\mathcal{I}_{+0}(x),\\ &0\leq x_{1,i}+d_{1,i}\perp x_{2,i}+d_{2,i}\geq 0,\forall i\in\mathcal{I}_{00}(x)\}.\end{split}$ (5)
Similarly to $x=(x_{0},x_{1},x_{2})$, the vector $d$ is partitioned as $d=(d_{0},d_{1},d_{2})\in\mathbb{R}^{p}\times\mathbb{R}^{m}\times\mathbb{R}^{m}$. Observe that at a feasible point $x\in\Omega$ all constraint residuals above are zero but they are kept for convenience. In contrast to the usual linearized feasible cone, here the combinatorial structure is kept for the degenerate index set ${\mathcal{I}}_{00}(x)$ and $\mathcal{F}_{\Omega}(x)$ is nonconvex, if $\mathcal{I}_{00}$ is nonempty.
In general, it holds that ${\mathcal{T}}_{\Omega}(x)\subseteq\mathcal{F}_{\Omega}(x)$ Flegel2005a . If the so-called MPEC-Abadie constraint qualification (MPEC-ACQ) is satisfied, i.e., ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$ Flegel2005a , we can use the $\mathcal{F}_{\Omega}(x)$ in Theorem 2.1. In particular, one can define a LPEC to characterize an algebraic B-stationary point (sometimes also called linearized B-stationarity Flegel2005a ), or B-stationary point for short.
Definition 2.1 (B-stationarity)
A point $x^{*}$ is called (algebraic) B-stationary if $d=0$ solves the following LPEC:
$\underset{d\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $\nabla f(x^{*})^{\top}d$ (6a)
s.t. $c(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,$ $\forall i\in{\mathcal{A}}(x^{*})$ (6b)
$x_{1,i}^{*}+d_{1,i}=0,\$$\forall i\in{\mathcal{I}}_{0+}(x^{*}),$ (6c)
$x_{2,i}^{*}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{*}),$ (6d)
$0\leq x_{1,i}^{*}+d_{1,i}\perp x_{2,i}^{*}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}).$ (6e)
To verify that $x^{*}$ is B-stationary, one has to verify that $d=0$ is a minimizer of an LPEC. Note that if $|{\mathcal{I}}_{00}(x^{*})|=0$, then (6) is an linear program (LP).
Next, we review multiplier-based stationary concepts for MPECs and some auxiliary NLPs associated with the MPEC (1). One of these problems will be solved by the algorithm introduced below, while others serve to define additional concepts relevant to MPECs. To this end, we introduce further index sets. Let $x^{*}\in\Omega$ and ${\mathcal{D}}_{1}(x^{*}),{\mathcal{D}}_{2}(x^{*})$ be a partition of the biactive set ${\mathcal{I}}_{00}(x^{*})$, with $\mathcal{D}_{1}(x^{*})\cup\mathcal{D}_{2}(x^{*})={\mathcal{I}}_{00}(x^{*})$ and $\mathcal{D}_{1}(x^{*})\cap\mathcal{D}_{2}(x^{*})=\emptyset$.
${\mathcal{I}}_{1}(x^{*})={\mathcal{I}}_{0+}(x^{*})\cup\mathcal{D}_{1}(x^{*}),\quad{\mathcal{I}}_{2}(x^{*})={\mathcal{I}}_{+0}(x^{*})\cup\mathcal{D}_{2}(x^{*}).$ (7)
Let us denote the set of all possible partitions by $\mathcal{P}(x^{*})=\{({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\}$. The number of possible partitions is $|\mathcal{P}(x^{*})|=2^{|{\mathcal{I}}_{00}(x^{*})|}$. For every partition, we can define a branch NLP, denoted by $\mathrm{BNLP}{({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))}$, as follows:
$\min_{x\in{\mathbb{R}}^{n}}\quad$ $f(x)$ (8a)
$c(x)\geq 0,$ (8b)
$x_{1,i}=0,\ x_{2,i}\geq 0,\;$ $\forall i\in{\mathcal{I}}_{1}(x^{*}),$ (8c)
$x_{1,i}\geq 0,\ x_{2,i}=0,\;$ $\forall i\in{\mathcal{I}}_{2}(x^{*}).$ (8d)
[…]
$\min_{x\in{\mathbb{R}}^{n}}\quad$ $f(x)$ (9a)
s.t. $c(x)\geq 0,$ (9b)
$x_{1,i}=0,\ x_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{0+}(x^{*}),$ (9c)
$x_{1,i}\geq 0,\ x_{2,i}=0,\;$ $\forall i\in\mathcal{I}_{+0}(x^{*}),$ (9d)
$x_{1,i}=0,\ x_{2,i}=0,\;$ $\forall i\in\mathcal{I}_{00}(x^{*}),$ (9e)
[…]
$x_{1,i}\geq 0,\ x_{2,i}\geq 0,\;\forall i\in\mathcal{I}_{00}(x^{*}).$
[…]
It can be seen that the following holds locally around a point $x^{*}\in\Omega$ Scheel2000 :
$\Omega_{\mathrm{TNLP}}=\bigcap_{(\mathcal{I}_{1},\mathcal{I}_{2}){\in\mathcal{P}}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}\subset\Omega=\bigcup_{(\mathcal{I}_{1},\mathcal{I}_{2}){\in\mathcal{P}}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1},\mathcal{I}_{2})}}.$ (10)
[…]
Definition 2.2
The MPEC (1) is said to satisfy the MPEC-LICQ (MPEC-MFCQ) at a feasible point $x^{*}\in\Omega$ if the corresponding $\mathrm{TNLP}$ (9) satisfies the LICQ (MFCQ) at the same point $x^{*}$.
Next, we define the MPEC Lagrangian. This is simply the standard Lagrangian for the BNLP/TNLP and reads as:
$\mathcal{L}^{\mathrm{{MPEC}}}(x,\mu,\nu,\xi)=f(x)-\mu^{\top}c(x)-\nu^{\top}x_{1}-\xi^{\top}x_{2},$ (11)
with the MPEC Lagrange multipliers $\mu\in{\mathbb{R}}^{n_{c}}$, $\nu\in{\mathbb{R}}^{m}$ and $\xi\in{\mathbb{R}}^{m}$.
By applying the KKT conditions to the TNLP (9), we obtain W-stationarity concept, and applying them to the RNLP yields S-stationary Scheel2000 . Further restrictions on the multipliers $\nu^{*}_{i},\xi^{*}_{i},\ i\in{\mathcal{I}}_{00}(x^{*})$, lead to M-, C-, and A-stationarity, which are not directly related to the KKT conditions of an auxiliary NLPs, see e.g.  Kim2020 ; Outrata1999 ; Scheel2000 . These concepts are summarized in the next definition.
Definition 2.3 (Stationarity concepts for MPECs)
For a feasible point $x^{*}\in\Omega$, we distinguish the following stationarity concepts.
[…]
    $\nabla_{x}\mathcal{L}^{\mathrm{{MPEC}}}(x^{*},\mu^{*},\nu^{*},\xi^{*})=0,$
    $0\leq\mu^{*}\perp c(x^{*})\geq 0,$
    $x^{*}_{1,i}\geq 0,\nu^{*}_{i}=0,\;$ $\forall i\in{\mathcal{I}}_{+0}(x^{*}),$
    $x^{*}_{2,i}\geq 0,\xi^{*}_{i}=0,\;$ $\forall i\in{\mathcal{I}}_{0+}(x^{*}),$
    $x^{*}_{1,i}=0,\;\nu_{i}^{*}\in{\mathbb{R}},\;$ $\forall i\in{\mathcal{I}}_{0+}(x^{*})\cup{\mathcal{I}}_{00}(x^{*}),$
    $x^{*}_{2,i}=0,\;\xi_{i}^{*}\in{\mathbb{R}},\;$ $\forall i\in{\mathcal{I}}_{+0}(x^{*})\cup{\mathcal{I}}_{00}(x^{*}).$
  • Strong stationarity (S-stationarity) Scheel2000 : A point $x^{*}$ is called S-stationary if it is weakly stationary and $\nu^{*}_{i}\geq 0,\xi^{*}_{i}\geq 0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
  • Clarke stationarity (C-stationarity) Scheel2000 : A point $x^{*}$ is called C-stationary if it is weakly stationary and $\nu^{*}_{i}\xi^{*}_{i}\geq 0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
  • Mordukhovich stationarity (M-stationarity) Outrata1999 : A point $x^{*}$ is called M-stationary if it is weakly stationary and if either $\nu^{*}_{i}>0$ and $\xi^{*}_{i}>0$ or $\nu^{*}_{i}\xi^{*}_{i}=0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
  • Abadie stationarity (A-stationarity) Flegel2005a : A point $x^{*}$ is called A-stationary if it is weakly stationary and $\nu^{*}_{i}\geq 0$ or $\xi^{*}_{i}\geq 0$ for all $i\in{\mathcal{I}}_{00}(x^{*})$.
Note that, if ${\mathcal{I}}_{00}=\emptyset$, then all stationarity concepts are identical and collapse to the concept of S-stationarity. From the definitions it follows that S-stationarity is the most restrictive since it has the tightest condition in the multipliers $\nu^{*}_{i},\xi^{*}_{i},i\in{\mathcal{I}}_{00}(x^{*})$. Weaker stationarity concepts, ordered by strength, are M, C, A, and W, with the stronger ones implying the weaker ones.
The relationship between multiplier-based stationarity (Def. 2.3) and B-stationarity (Def. 2.1) is established in the following theorem.
Theorem 2.2
((Scheel2000, , Theorem 4)) If $x^{*}$ is an S-stationary point of the MPEC (1), then it is also B-stationary. If, in addition, the MPEC-LICQ holds, then every B-stationary point is S-stationary.
Remark 2.1
In summary, without assuming MPEC-LICQ, B-stationarity cannot be verified using multiplier-based concepts. This assumption is essential, as the next weaker condition, MPEC-MFCQ, does not guarantee S-stationarity (cf. (Scheel2000, , Example 3)). Weaker termination criteria than B-stationarity may be considered unsatisfactory, since even M-stationarity allows first-order descent directions Leyffer2007 . Even under MPEC-LICQ, regularization Ralph2004 ; Hoheisel2013 ; Kanzow2015 and some active set methods (Kirches2022, , Section 3) may converge to points weaker than S-stationary. Therefore, we propose a method using the LPEC (6), to certify B-stationarity.
3 MPECopt: A globally convergent method for computing B-stationary points of MPECs
This section presents the MPEC Optimizer (MPECopt), an easy-to-implement active set method that solves a finite sequence of LPECs or LPs, and BNLPs. It returns either a B-stationary point or a stationary point of the problem of minimizing the constraint infeasibility. The latter is a common outcome of nonlinear programming algorithms Fletcher2002 ; Waechter2006 .
Before providing all the details, we summarize the main ideas behind the new algorithm in Section 3.1. Thereafter, Section 3.2 discusses different formulations of LPECs, their properties, and how LPECs can be solved numerically. MPECopt consists of two phases. Details of Phase I and Phase II are discussed in Sections 3.3 and  3.4, respectively.
3.1 Summary of the approach
Algorithm 1 MPECopt: Computing a B-stationary point of MPECs
1: Call Phase I to find a feasible point $x^{0}$ of the MPEC (1); set $k=0$
2: while $d^{k,l}\neq 0$ do $\triangleright$ major iterations
3:   $x^{k,0}=x^{k}$
4:   for $l=0,\ldots,$ do $\triangleright$ minor iterations
5:    Solve $\textrm{LPEC}(x^{k},\rho^{k,l})$ in (15) to obtain $d^{k,l}$
6:    if $d^{k,l}=0$ then terminate $\triangleright$ B-stat. point found
7:    Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$ to obtain $x^{k,l}$
8:    if BNLP unbounded then terminate $\triangleright$ MPEC is unbounded
9:    else if $f(x^{k,l})<f(x^{k})$ then
10:      Set $x^{k+1}\!=\!x^{k,l}$, k=k+1k\!=\!k\!+\!1, increase trust-region radius $\rho^{k+1,0}$, and break for-loop
11:    else
12:      Reduce trust-region radius $\rho^{k,l+1}$
Algorithm 1 provides simplified pseudo-code for MPECopt, which consists of two phases. Phase I either finds an MPEC-feasible point $x^{0}\in\Omega$ or a stationary point of the problem of minimizing the constraint infeasibility. Phase II is summarized in lines 2–11 of Algorithm 1, while the details of our actual implementation are given in Section 3.4. The outcome of Phase II is that it finds a B-stationary point of the MPEC or detects that the problem is unbounded.
If a feasible point is found in Phase I, Phase II creates a finite sequence of iterates $x^{k}\in\Omega$ by solving BNLPs and LPECs. If Phase II finds that any BNLP is unbounded, then the MPEC is unbounded, and the algorithm terminates. The LPECs are similar to (6), but they are defined at the linearization point $x^{k}$ and include an additional trust-region constraint $\|d\|_{\infty}\leq\rho$, $\rho>0$. These LPECs are compactly denoted by $\mathrm{LPEC}(x,\rho)$; see Section 3.2 for details.
At each major iteration $k$, an inner loop with counter $l$ solves $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$ and, possibly, a corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}(x^{k}+d^{k,l}))$, until either a feasible point $x^{k+1}=x^{k,l}$ with an improved objective value is found (lines 7–10), or the algorithm terminates due to $d^{k,l}=0$ (line 6) or the MPEC is found to be unbounded (line 8).
[…]
Note that the LPEC steps $x^{k}+d^{k,l}$ are never used to update iterates directly. This avoids cases where, under nonlinear constraints, $x^{k}+d^{k,l}\notin\Omega$ while $f(x^{k}+d^{k,l})$ increases, which would require sophisticated globalization for the LPEC steps. Instead, we accept only BNLP stationary points $x^{k,l}$, which remain feasible for the MPEC (1).
3.2 The LPEC subproblems
In Section 3.2.1 we discuss and compare two different formulations for the LPEC used in our method, and in Section 3.2.2 we discuss how they are efficiently solved.
3.2.1 On different LPEC formulations
The first formulation in inspired by Def. 2.1. Given a feasible point $x^{k}\in\Omega$, Phase II in Algorithm 1 requires solving the following $\mathrm{LPEC}(x^{k},\rho)$:
$\underset{d\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $\nabla f(x^{k})^{\top}d$ (12a)
s.t. $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,$ ${\forall i\in{\mathcal{A}}(x^{k})}$ (12b)
$x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{0+}(x^{k}),$ (12c)
$x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{k}),$ (12d)
$0\leq x_{1,i}^{k}+d_{1,i}\perp x_{2,i}^{k}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{00}(x^{k}),$ (12e)
$\|d\|_{\infty}\leq\rho.$ (12f)
Compared to (6) in Def. 2.1, the LPEC above has the following modifications. Additional to (6c) and (6d), here we include the constraint $x_{2,i}^{k}+d_{2,i}\geq 0$ and $x_{1,i}^{k}+d_{1,i}\geq 0$ in (12c) and (12d), respectively. This ensures that for any solution $d$ of (12) the point $x^{k}+d$ satisfies the complementarity constraints, which is necessary for a meaningful definition of the active sets in  (7). Moreover, we add a trust region constraint for the step $d$ in (12f), where $\rho>0$ is the trust region radius. This constraint guarantees that the LPEC is bounded. We use an $\ell_{\infty}$ norm for the step, yielding simple lower and upper bounds on $d$.
If $d=0$ is a global minimizer of (12), then none of the additionally introduced constraints is active, and moreover, if the MPEC-ACQ holds, then by Theorem 2.1, $x^{k}$ is a B-stationary point of MPEC (1). On the other hand, if $x^{k}$ is not B-stationary (i.e., $d=0$ is not a minimizer of (12)), and if ${\mathcal{A}}(x^{k}+d)={\mathcal{A}}(x^{k})$, then $\nabla f(x^{k})^{\top}d<0$ and $d$ is a first-order descent direction. In this case, we take the sets ${\mathcal{I}}_{1}(x^{k}+d)$ and ${\mathcal{I}}_{2}(x^{k}+d)$ as the new active set guess for the next BNLP to be solved in Phase II; more details on the actual selection of these index sets are provided below.
Next, since algebraic B-stationarity is stronger than geometric B-stationarity, we show that the assumption of MPEC-ACQ (i.e., that ${\mathcal{T}}_{\Omega}(x)=\mathcal{F}_{\Omega}(x)$), cannot be relaxed if one wishes to verify geometric B-stationarity (cf. Theorem 2.1) by solving the LPEC (12), which verifies algebraic B-stationarity (cf. Def. 2.1).
(a) The MPEC (13).
(b) LPEC (14) $\rho=1$.
(c) LPEC (14) with $\rho=0.5$.
Figure 1: Illustration of Example 3.1. If a constraint qualification does not hold, the LPEC cannot verify geometric B-stationary.
Example 3.1
Consider the following MPEC:
[…]
The point $x^{*}=(0,0)$ is the global optimum of this problem and thus geometric B-stationary. It can be seen that the tangent cone of this problem is ${\mathcal{T}}_{\Omega}(x^{*})=\{d\in{\mathbb{R}}^{2}\mid d_{1}=0,d_{2}\geq 0\}$. The corresponding LPEC reads as:
$\underset{d\in{\mathbb{R}}^{2}}{\mathrm{min}}\;\quad$ $-2d_{1}+d_{2}$ (14a)
s.t. $d_{2}\geq 0$ (14b)
$0\leq d_{1}\perp d_{2}\geq 0$ (14c)
$\|d\|_{\infty}\leq\rho.$ (14d)
Clearly, ${\mathcal{T}}_{\Omega}(x^{*})\subset\mathcal{F}_{\Omega}(x^{*})$, which means that the MPEC-ACQ does not hold at $x^{*}$. The only solution of (14) is $d=(\rho,0)$. Therefore, this LPEC cannot verify geometric B-stationarity for any $\rho>0$. Fig. 1 illustrates the feasible sets of the MPEC and the LPEC for different $\rho$.
The LPEC (12) is well-defined at feasible points. However, in the Phase I algorithm of Section 3.3, we must also consider LPECs at infeasible points, where index sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+},{\mathcal{I}}_{00}$ are note defined. Existing strategies for identifying active sets from infeasible points often require overly restrictive assumptions Lin2006 . Inspired by (Scholtes2004, , Sec. 3.2), alternative approaches instead include all linearized constraints, yielding the LPEC Kirches2022 ; Leyffer2007 :
$\underset{d\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $\nabla f(x^{k})^{\top}d$ (15a)
s.t. $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,$ $\forall i\in\{1,\ldots,n_{c}\},$ (15b)
$0\leq x_{1,i}^{k}+d_{1,i}\perp x_{2,i}^{k}+d_{2,i}\geq 0,$ $\forall i\in\{1,\ldots,m\},$ (15c)
$\|d\|_{\infty}\leq\rho.$ (15d)
We refer to the LPEC (12) as the “reduced” LPEC and (15) as the “full” LPEC. For feasible points $x^{k}$, the full LPEC contains the feasible set of the reduced one. Moreover, since $d=0$ is always feasible for $x^{k}\in\Omega$, the optimal objective values of both (12) and (15) are nonpositive, regardless of the trust region radius $\rho>0$. At infeasible points $x^{k}\notin\Omega$, the optimal objective value of (15) can take any value in $\mathbb{R}$}. Moreover, at such points, the trust-region radius $\rho$ must be large enough to ensure that the LPEC is feasible, cf. Theorem 4.1.
Let us introduce the following notation shorthand ${\mathcal{I}}_{+0}^{k}:={\mathcal{I}}_{+0}(x^{k})$ and ${\mathcal{I}}_{0+}^{k}:={\mathcal{I}}_{0+}(x^{k})$. Denote the set of inactive inequality constraints by $\bar{{\mathcal{A}}}^{k}:=\bar{{\mathcal{A}}}(x^{k})=\{i\in\{1,\ldots,n\}\mid c_{i}(x)>0\}$. For a feasible point $x^{k}\in\Omega$, the solution sets of the two LPECs are related as follows. The proof can be found in Appendix A.
Lemma 3.1
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1). For all trust region radii that satisfy
$\begin{split}0<\rho<\bar{\rho}=\min\Big\{&\{{x^{k}_{1,i}}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{{x^{k}_{2,i}}\mid i\in{\mathcal{I}}_{0+}^{k}\}\\ &\cup\{{c_{i}(x^{k})}\|\nabla c_{i}(x^{k})\|^{-1}\mid i\in\bar{{\mathcal{A}}}^{k},\|\nabla c_{i}(x^{k})\|\neq 0\}\Big\},\end{split}$ (16)
the sets of the local minimizers of  (12) and (15) are identical. In the special case of ${\mathcal{I}}_{00}^{k}=\{1,\ldots,m\}$ and $\bar{{\mathcal{A}}}^{k}=\emptyset$ the reduced and full LPEC coincide and in that case, $\bar{\rho}=\infty$.
In other words, at a feasible point $x^{k}$ and for a sufficiently small trust-region radius, all branches that are not biactive at $x^{k}$ are eliminated, and the linearizations of inactive constraints remain inactive. Thus, the feasible sets of the two LPECs coincide. On the other hand, for $\rho\geq\bar{\rho}$, the global minimizer $d=0$ of (12) may correspond only to a local minimizer of (15), cf. Example 3.2. Hence, global optimality of $d=0$ is sufficient but not necessary to verify B-stationarity via (15Scholtes2004 .
In the next example, we show one more advantage and disadvantage of the full over the reduced LPEC.
(a) $\rho=0.6$.
(b) $\rho=1.2$.
(c) $\rho=1.2$.
Figure 2: Illustration of the MPEC  (17). The yellow stars denote the two stationary points. The figure illustrates the LPEC’s (15) trust region, objective gradient, and optimal solution $d$ for different values of $\rho$ at different stationary points.
Example 3.2 (Finding a better objective value)
Consider the following MPEC:
$\min_{x\in{\mathbb{R}}^{2}}\ 4(x_{1}-1)^{2}+(x_{2}-1)^{2}\quad\textnormal{s.t.}\quad 0\leq x_{1}\perp x_{2}\geq 0,$ (17)
The example has two S-stationary points, $\bar{x}=(1,0)$ and $\hat{x}=(0,1)$, both with empty biactive sets, and respective objective values $f(\bar{x})=1$ and $f(\hat{x})=4$. In Fig. 2a, if $\rho$ is sufficiently small or the reduced LPEC is used, the point $\hat{x}$ with $f(\hat{x})=4$ is verified as B-stationary point. On the other hand, with a larger trust region radius, as depicted in Fig. 2b, the globally optimal solution of the full LPEC identifies a BNLP with a lower objective value $f(\bar{x})=1$. Conversely, if we start at $\bar{x}$, the full LPEC can identify a BNLP with a larger objective, as depicted in Fig. 2c. However, this step will be rejected, and the trust region radius reduced until $\bar{x}$ is verified as B-stationary.
[…]
Lemma 3.2
Let $x^{k}\in\Omega$ be a feasible point of the MPEC (1) and suppose that the MPEC-MFCQ holds at this point. Let $d$ be any feasible point of the $\mathrm{LPEC}(x^{k},\rho)$ in (15) for $\rho\in(0,\bar{\rho})$, where $\bar{\rho}$ is defined in (16). Then $x^{k}$ is a feasible point of $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}+d),{\mathcal{I}}_{2}(x^{k}+d))$.
A proof is given in Appendix A.
3.2.2 On solving LPECs
LPECs are nonconvex and combinatorial problems, and the effectiveness of our proposed method relies on solving them efficiently. In practice, our computational results indicate that although the worst-case complexity of LPECs may be exponential, we typically only need to solve the root node of the MILP reformulation described below.
We summarize several favorable cases that frequently arise. First, if the LPEC contains only bound and complementarity constraints, it can be solved in linear time (Kirches2022, , Propositions 2.3 and 2.4). This algorithm is incorporated into our implementation. Second, at a feasible point $x^{k}$, if ${\mathcal{I}}_{00}(x^{k})=\emptyset$, then the LPEC (12) reduces to a standard LP. According to (10), the LPEC can then be solved by enumerating all branch LPs, which is not computationally expensive when $|{\mathcal{I}}_{00}(x^{k})|$ is small. Third, if $x^{k}$ is not B-stationary, it is unnecessary to solve all branch LPs; it suffices to find a feasible direction $d$ with $\nabla f(x^{k})^{\top}d<0$ (cf. Corollary 4.2). Moreover, in the LPECs arising in Phase I of our method, it is sufficient to simply find a feasible point (cf. Theorem 4.1). In a MILP formulation, the solver can therefore be terminated as soon as such a feasible point is obtained. Section 5 shows that this results in a significant reduction in the number of LPs solved in the MILP algorithm. Finally, if $x^{k}$ is S-stationary, B-stationarity can be verified by solving a single LP (cf. Proposition 3.1). In summary, except when verifying B-stationarity, it is rarely necessary to solve the LPEC to global optimality – and even then, the task often reduces to solving a single LP.
In this paper, we focus on arguably the easiest-to-implement approach for solving LPECs, which does not suffer from the drawbacks of regularization methods. We consider the equivalent MILP formulation of LPEC (15Hu2012 :
$\underset{{d\in{\mathbb{R}}^{n},\ y\in\{0,1\}^{m}}}{\mathrm{min}}\;\quad$ $\nabla f(x^{k})^{\top}d$ (18a)
s.t. $c(x^{k})+\nabla c(x^{k})^{\top}d\geq 0,$ (18b)
$0\leq x^{k}_{1,i}\!+\!d_{1,i}\!\leq\!{y_{i}}M,$ $\forall i\in\{1,\ldots,m\},$ (18c)
$0\leq x^{k}_{2,i}\!+\!d_{2,i}\!\leq\!{(1-y_{i})}M,$ $\forall i\in\{1,\ldots,m\},$ (18d)
$\|d\|_{\infty}\leq\rho.$ (18e)
The constraints (18c) and (18d) correspond to the big-M reformulation of the complementarity constraints in (15c), where $M$ is a sufficiently large positive constant. The binary variables $y\in\{0,1\}^{m}$ effectively select one of the branches of the L-shaped set defined by the complementarity constraints.
To reduce the problem complexity for a given $\rho$, if $d_{i}\leq\rho<x_{1,i}$ or $d_{i}\leq\rho<x_{2,i}$, we can fix the corresponding binary variable $y_{i}$ for every $i$ in a preprocessing step, as the other branch becomes unreachable. In the special case of selecting $\rho$ as in (16), all binary variables $y_{i}$ for $i\notin{\mathcal{I}}_{00}(x^{k})$ can be determined in advance, and it is observed that this contributes to the computational efficiency of our approach.
The MILP (18) can be solved with state-of-the-art commercial and open-source solvers such as Gurobi Gurobi or HiGHS Huangfu2018 . We discuss some special properties of the MILP (18), which are favorable for MILP solvers Gurobi ; Huangfu2018 . We denote the optimal objective value of (18) by $V_{\mathrm{MILP}}(x^{k})$, and of its LP relaxation, that is, a problem where $y\in\{0,1\}^{m}$ is replaced by $y\in[0,1]^{m}$, by $V_{\mathrm{LP}}(x^{k})$. Thus, it holds that $V_{\mathrm{LP}}(x^{k})\leq V_{\mathrm{MILP}}(x^{k})$. The tighter the relaxation, i.e., the smaller $|V_{\mathrm{MILP}}(x^{k})-V_{\mathrm{LP}}(x^{k})|$ is, the less branching may be needed in MILP solvers Wolsey1998 . Big-M formulations can be inefficient because they can lead to rather loose relaxations of the MILP Hu2008 . However, for $\rho<\frac{M}{2}$, the trust-region constraint (18e) always dominates the bounds on $d$ from the big-M constraints (18c) and (18d). Thus, $M$ only needs to be large enough to ensure the feasibility of (18c)–(18d). A suitable choice is $M=\max(x_{1}^{k},x_{2}^{k})+\rho$. In addition, observe that the constraint (18e) makes the feasible set of the MILP compact. Moreover, for feasible points $x^{k}$, $\rho$ can be arbitrarily small, leading to a smaller $V_{\mathrm{LP}}(x^{k})$ and thus tighter relaxations of the MILP.
Furthermore, given a feasible complementarity pair, e.g. from the previous BNLP, a feasible binary vector $y$ can be constructed, providing an upper bound for $V_{\mathrm{MILP}}(x^{k})$. The availability of an upper bound can reduce branching in branch-and-bound MILP algorithms Wolsey1998 , and our implementation includes the procedure for computing such a feasible binary vector. Our numerical experiments indicate that these factors typically result in not large numbers of LP that need to be solved when solving the MILP (18). Interestingly, in our experiments the MILP formulations for LPECs are often substantially faster than solving the LPEC (15) with noncombinatorial regularization-based methods, cf. Appendix C.
We discuss a special case that frequently arises in our numerical experiments, namely when $x^{*}\in\Omega$ is S-stationary. In this case, the MILPs are particularly easy to solve since the relaxation is tight, i.e., $V_{\mathrm{MILP}}(x^{*})=V_{\mathrm{LP}}(x^{*})$. Thus, verifying B-stationarity reduces to solving a single LP in the MILP solver. We prove this for the reduced LPEC (12). However, if $\rho<\bar{\rho}$ (with $\bar{\rho}$ from (16)) and the preprocessing step for fixing integers is applied, then the feasible sets of the full MILP (18) and reduced MILP (19) coincide.
Proposition 3.1
Let $x^{*}\in\Omega$ be an S-stationary point of the MPEC (1) at which the MPEC-MFCQ holds. Consider the MILP reformulation of the reduced LPEC (12):
$\underset{d\in{\mathbb{R}}^{n}\!,y\in\{\!0,1\!\}^{|{\mathcal{I}}_{00}^{*}|}}{\mathrm{min}}\;\quad$ $\nabla f(x^{*})^{\top}d$ (19a)
s.t. $c_{i}(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,$ $\forall i\in{\mathcal{A}}(x^{*}),$ (19b)
$x_{1,i}^{*}+d_{1,i}=0,\ x_{2,i}^{*}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{0+}(x^{*}),$ (19c)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{*}),$ (19d)
$0\leq x^{*}_{1,i}\!+\!d_{1,i}\!\leq\!{y_{i}}M,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}),$ (19e)
$0\leq x^{*}_{2,i}\!+\!d_{2,i}\!\leq\!{(1-y_{i})}M,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}),$ (19f)
$\|d\|_{\infty}\leq\rho.$ (19g)
For $\rho<\frac{M}{2}$, $d=0$ is a global minimizer of both the MILP (19) and its relaxation with $y_{i}\in[0,1]$.
Proof. First we show that $d=0$ is global minimizer of the MILP (19). It follows from Theorem 2.2 that if $x^{*}\in\Omega$ is S-stationary it is also B-stationary. Hence, $d=0$ is a global minimizer of the LPEC (6) and of ${\mathrm{LPEC}}(x^{*},\rho)$ in (12), because the additional trust-region constraint in the latter remains inactive. The MILP (19) and LPEC (12), are equivalent, hence $d=0$ is a global minimizer of the MILP.
[…]
$\underset{{d\in{\mathbb{R}}^{n}}}{\mathrm{min}}\;\quad$ $\nabla f(x^{*})^{\top}d$ (20a)
s.t. $c_{i}(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,\ \forall i\in{\mathcal{A}}(x^{*}),$ (20b)
$x_{1,i}^{*}+d_{1,i}=0,\ x_{2,i}^{*}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{0+}(x^{*}),$ (20c)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}=0,\;$ $\forall i\in\mathcal{I}_{+0}(x^{*}),$ (20d)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{00}(x^{*}),$ (20e)
$\|d\|_{\infty}\leq\rho$ (20f)
[…]
In the relaxed MILP, $y_{i}\in\{0,1\}$ in (19) is replaced by $y_{i}\in[0,1]$. Next, from the upper bounds in (19e) and (19f), we have that $\frac{d_{1,i}}{M}\leq y_{i}\leq 1-\frac{d_{2,i}}{M}$. Since the MILP is feasible, it follows that this interval is nonempty if $\frac{d_{1,i}}{M}\leq 1-\frac{d_{2,i}}{M})$ holds. This implicitly defines the constraint: $d_{1,i}+d_{2,i}\leq M$. If we choose $\rho<\tfrac{M}{2}$, then the upper bounds in (19e) and (19f) are never active. For such a choice, the feasible sets of the relaxed MILP (19) and the LP (20) coincide for all components $d\in{\mathbb{R}}^{n}$, for all $y\in[0,1]^{|{\mathcal{I}}_{00}|}$. Since the objective of (19) does not depend on $y$, it follows that $d=0$ is also a global minimizer of the relaxation of (19). In conclusion, $d=0$ is a global solution of the MILP (19) and LPEC (12) (verifies B-stationarity), relaxed LP (verifies S-stationarity) and relaxed MILP. ∎
Remark 3.1 (On Algorithm 1 and mixed-integer algorithms)
Any MPEC (1) can be reformulated into a mixed-integer NLP (MINLP) using the big-M reformulation:
$\underset{x\in{\mathbb{R}}^{n},y\in\{0,1\}^{m}}{\mathrm{min}}\;\quad$ $f(x)$ (21a)
s.t. $c(x)\geq 0,$ (21b)
$0\leq x_{1,i}\leq y_{i}M,$ $\forall i\in\{1,\ldots,m\},$ (21c)
$0\leq x_{2,i}\leq(1-y_{i})M,$ $\forall i\in\{1,\ldots,m\}.$ (21d)
However, the convergence properties our method and MINLP methods differ in several aspects. Most MINLP methods converge to the global optimum when the relaxed integer problem is convex Belotti2013 , but Algorithm 1 does not. This distinction is illustrated by the MPEC example in Example 3.2, whose equivalent MINLP (after relaxing $y\in\{0,1\}^{m}$ to $y\in[0,1]^{m}$) yields a convex quadratic program. There, Algorithm 1 converges only to a B-stationary point, not necessarily the global optimum.
While Algorithm 1 solves NLPs and MILPs similar to many MINLP methods Belotti2013 , there are significant differences in termination criteria and the particular subproblems that are solved. Nonlinear branch-and-bound algorithms terminate when the optimality gap falls below a threshold Belotti2013 , which does not directly relate to B-stationarity unless the global optimum is found. Outer approximation algorithms also solve sequences of NLPs and MILPs Belotti2013 , but their MILPs include linearized constraints from multiple different linearization points, making them different from the MILP (18). Consequently, it is not immediately clear when these methods terminate at B-stationary points. In summary, Algorithm 1 is not a MINLP algorithm; rather, it employs MILP solvers as a tool for solving LPECs. It aims to find a locally stationary point of an MPEC, which—unlike in MINLPs – is a well-defined notion.
3.3 Phase I: Finding a feasible branch NLP
In this section, we introduce an algorithm to find an initial feasible BNLP and its corresponding solution $x^{0}$, or return a stationary point of an auxiliary problem that minimizes a measure of constraint infeasibility.
Regularization-based methods may not always find B-stationary points, but they can be efficient in finding (almost) feasible points. We outline a Phase I method that combines solving LPECs with regularization-based methods. While any regularization-based method can be used in this framework, we focus in our exposition on Scholtes’ global relaxation method, which performs well in numerical experiments Kim2020 ; Hoheisel2013 ; Nurkanovic2024b . It solves a sequence of NLPs of the form:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}$ $f(x)$ (Reg($\tau$))
s.t. $c(x)\geq 0,$
$x_{1},x_{2}\geq 0,$
$x_{1,i}x_{2,i}\leq\tau,$ $\forall i\in\{1,\ldots,m\}.$
We compute $x^{*}(\tau)$, a stationarity point of Reg($\tau$), which is not necessarily feasible for the MPEC (1). An LPEC is then constructed at $x^{*}(\tau)$ (for some $\tau>0$), and its solution determines an active set corresponding to a feasible BNLP.
It is well-known that certifying infeasibility of a general nonconvex optimization problem is difficult. In many NLP solver implementations, if the globalization strategy cannot make sufficient progress toward a feasible point, a feasibility restoration phase is triggered, cf. (Fletcher2002, , Sec. 3.3.), (Waechter2006, , Sec 3.3.), and (Byrd2006, , Sec. 6). In this step, an auxiliary NLP that minimizes the constraint violation is solved Byrd2006 ; Fletcher2002 ; Waechter2006 . If its stationary point is infeasible for the original problem, the solver terminates. Infeasibility of the original problem is sufficient but not necessary for this outcome. This outcome is often called local infeasibility Waechter2006 , and it is not a global certificate of infeasibility.
Lemma 3.3
Regard the MPEC (1) and the corresponding $\mathrm{Reg}(\tau)$ in  Reg($\tau$) for some $\tau>0$. If $\mathrm{Reg}(\tau)$ is infeasible then the same holds for the MPEC (1).
Proof. This follows from the fact that the feasible set of $\mathrm{Reg}(\tau)$ is a superset of $\Omega$. ∎
On the other hand, if Reg($\tau$) is feasible for all $\tau$, then also the MPEC (1) is feasible as well. We expect that if $\tau^{k}$ is sufficiently small, then the feasible set of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho)$ is locally a good approximation of the original feasible set, and can help identifying a feasible BNLP. Algorithm 2 details such a procedure. Before going into details, we illustrate how an LPEC solution may fail or succeed in identifying a feasible BNLP for different values of $\tau$.
Example 3.3
Consider the following MPEC:
[…]
with $a>1$. In the example, we set $a=1.1$. For larger values of $\tau$ both branches are feasible for the LPEC (15), but only the branch $x_{1}=0,x_{2}\geq 0$ is feasible for the MPEC. On the other hand, for smaller values of $\tau$, also in the LPEC, only the branch $d_{1}=0,\ d_{2}\geq 0$ is feasible, thus the LPEC identifies only active sets $({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$, which results in a feasible BNLP. Fig. 3 illustrates two sample solutions for $\mathrm{Reg}(\tau)$ (left plots) and the corresponding $\mathrm{LPEC}(x^{*}(\tau),\rho)$ (right plots). In this example, the smaller the parameter $a$ is the smaller $\tau$ must become.
(a) $\tau=1.0$.
(b) $\tau=1.0$.
(c) $\tau=0.04$.
(d) $\tau=0.04$.
Figure 3: Illustrating of Example 3.3, where the LPEC (15) can correctly and wrongly identify a feasible BNLP depending on the linearization point $x^{*}(\tau)$. Top: the linearization point $x^{*}(\tau)$ is not close enough to the feasible set, so the LPEC admits an MPEC-infeasible branch. Bottom: $x^{*}(\tau)$ sufficiently close, so all LPEC branches are MPEC-feasible.
Algorithm 2 Regularization-based Phase I of MPECopt
1:Input: $\rho_{I}>0$, $\kappa\in(0,1)$, $\tau^{0}>0$, $N^{\tau}>1$
2: while do
3:   Solve $\mathrm{Reg}(\tau^{k})$ for $x^{*}(\tau^{k})$
4:   if $\mathrm{Reg}(\tau^{k})$ locally infeasible then
5:    break $\triangleright$ MPEC locally infeasible
6:   else if $x^{*}(\tau^{k})\in\Omega$ then
7:    break; return $x^{0}=x^{*}(\tau^{k})$ $\triangleright$ $\mathrm{Reg}(\tau^{k})$ solution feasible for MPEC
8:   else if $h_{\perp}(x(\tau^{k}))<\rho_{I}$ then
9:    Solve $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho_{I})$ for $d^{k}$
10:    if $d^{k}$ exists then
11:      Solve $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ for $x^{0}$.
12:      if Solution exists then
13:       break, return $x^{0}$ $\triangleright$ feasible point found
14:   else
15:    $\tau^{k+1}=\kappa\tau^{k}$. $\triangleright$ no feasible LPEC/BNLP found, reduce homotopy parameter
We show below in Corollary 4.1 that if the MPEC admits a feasible point, then for sufficiently small $\tau$, then a feasible point of the LPEC yields a complementarity active set corresponding to a feasible BNLP. Consequently, the while loop in Algorithm 2 has finite termination. In a practical implementation, a point is declared feasible if the constraint residual is below a certain tolerance tolh. Formally, we require $h(x)\leq\texttt{tol}_{h}$, where $h(x)$ is the total infeasibility defined as:
$\begin{split}&h(x)=\max(h_{c}(x),h_{\!\perp\!}(x)),\\ &h_{c}(x)=\max(c^{-}(x)),\ \!h_{\!\perp\!}(x)\!=\!\max(|\min(x_{1},x_{2})|),\end{split}$ (22)
with $c^{-}(x)=\min(0,c(x))$.
If $\mathrm{Reg}(\tau^{k})$ is infeasible (line 5), or its solution is feasible for the MPEC (line 7), the algorithm terminates. Otherwise, the solution $x^{*}(\tau^{k})$ is used to construct ${\mathrm{LPEC}}(x^{*}(\tau^{k}),\rho_{I})$. The fixed trust-region radius $\rho_{I}>0$ must be sufficiently large to ensure feasibility of this LPEC. Since a solution of $\mathrm{Reg}(\tau^{k})$ already satisfies the general inequality constraints, i.e., $c(x^{*}(\tau^{k}))\geq 0$, the smallest value of $\rho_{I}$ required for feasibility is $\rho_{I}>h_{\perp}(x^{*}(\tau^{k}))$. The LPEC is solved only if this inequality holds (line 8). If it is not yet satisfied or the LPEC is infeasible, the homotopy parameter is reduced to obtain a better linearization point $x^{*}(\tau^{k})$ (line 15). If a solution to the LPEC is found, Algorithm 2 attempts to solve the corresponding $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ (line 11). If this is successful, a feasible point is obtained. Otherwise, the homotopy parameter is reduced to compute a better linearization point $x^{*}(\tau^{k+1})$.
Theorem 4.1 shows that if $x^{*}(\tau^{k})$ is sufficiently close to a feasible point $x\in\Omega$, then $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau^{k})+d^{k}),{\mathcal{I}}_{2}(x^{*}(\tau^{k})+d^{k}))$ is feasible, where $d^{k}$ is a feasible point of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho^{I})$. Moreover, under MPEC-MFCQ, the sequence $x^{*}(\tau^{k})$ converges to a stationary and thus feasible point of the MPEC Hoheisel2013 ; Kanzow2015 .
Remark 3.2
Algorithm 2 with the outcomes in lines 7 or 13 can also be viewed as a standalone method that computes a stationary point of MPEC (1c), though not necessarily a B-stationary point. However, if the computed point is B-stationary, this will be immediately verified in Phase II. If the trust-region radius is sufficiently small (cf. Lemma 3.1), the first LPEC to be solved will have a solution $d=0$. This behavior was occasionally observed in our numerical experiments in Section 5.
3.4 Phase II: Computing a B-stationary point
Algorithm 3 provides a detailed version of the Phase II algorithm outlined in Algorithm 1. It consists of two nested loops: an outer loop with counter $k$, which makes progress toward B-stationary points, and an inner loop with counter $l$, which controls the trust-region radius $\rho^{k,l}$ and solves LPECs and BNLPs until a point $x^{k,l}\in\Omega$ with improved objective value is found, i.e., $f(x^{k,l})<f(x^{k})$. Under reasonable assumptions, the inner loop is guaranteed to terminate in finitely many steps; cf. Proposition 4.3. With additional conditions, the outer loop may also terminate finitely; cf. Theorem 4.2.
Algorithm 3 Phase II of MPECopt
1:Input: $x^{0}\in\Omega$, $\gamma^{\mathrm{L}}\in(0,1)$, $N^{\mathrm{out}}$, $N^{\mathrm{in}}$, $\rho^{\mathrm{lb}}$, $\rho^{\mathrm{ub}}$.
2: for $k=0,\ldots,N^{\mathrm{out}}$ do $\triangleright$ major/outer iterations
3:   Reset trust region radius $\rho^{k,0}\in[\rho^{\mathrm{lb}},\rho^{\mathrm{ub}}]$;
4:   for $l=0,\ldots,N^{\mathrm{in}}$ do $\triangleright$ minor/inner iterations
5:    Solve $\mathrm{LPEC}(x^{k},\rho^{k,l})$ for $d^{k,l}$
6:    if $d^{k,l}=0$ then
7:      terminate $\triangleright$ B-stationary point found
8:    else
9:      Select partition $({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))$ via Eq. (23)
10:      if $l=0$ or ${\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l})\neq{\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l-1})$ then
11:       Solve $\mathrm{BNLP}({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))$ to obtain $x^{k,l}$
12:       if BNLP solver failed or $f(x^{k,l})\geq f(x^{k})$ then
13:         ${\rho}^{k,l+1}={\gamma^{\mathrm{L}}\rho^{k,l}}$; $\triangleright$ reduce TR radius
14:       else
15:         $x^{k+1}=x^{k,l}$; break inner for-loop; $\triangleright$ step accepted
16:      else
17:       ${\rho}^{k,l+1}={\gamma^{\mathrm{L}}\rho^{k,l}}$ $\triangleright$ reduce TR radius
Algorithm 3 starts from a feasible point $x^{0}\in\Omega$. At the beginning of each outer iteration, the trust-region radius is reset to $\rho^{k,0}\in[\rho^{\mathrm{lb}},\rho^{\mathrm{ub}}]$ with $\rho^{\mathrm{ub}}>\rho^{\mathrm{lb}}>0$. In theory any initial $\rho^{k,0}$ from the interval can be picked. In our implementation, we set $\rho^{k,0}=10^{-3}$, with $\rho^{\mathrm{ub}}=10^{3}$, and $\rho^{\mathrm{lb}}=10^{-7}$.
[…]
If $d^{k,l}\neq 0$, the algorithm proceeds by selecting and solving the corresponding BNLP (lines 911). If ${\mathcal{I}}_{00}(x^{k}+d^{k,l})\neq\emptyset$, the choice of the next partition is not unique. Proposition 4.3 ensures that there exists a threshold $\bar{\rho}$ such that for all $\rho^{k,l}<\bar{\rho}$, every possible partition from $\mathcal{P}(x^{k}+d^{k,l})$ yields a BNLP whose local minimizer improves the objective. Therefore, we may pick any partition from this set. We denote the selected partition by $({\mathcal{I}}_{1}^{*}(x^{k}+d^{k,l}),{\mathcal{I}}_{2}^{*}(x^{k}+d^{k,l}))\in\mathcal{P}(x^{k}+d^{k,l})$. In our implementation, we use the MILP solution to pick the partition:
${\mathcal{I}}^{*}_{1}(x^{k}+d^{k,l})=\{i\mid y_{i}^{k,l}=0\},\ {\mathcal{I}}^{*}_{2}(x^{k}+d^{k,l})=\{i\mid y_{i}^{k,l}=1\}.$ (23)
Here the integer values $y_{i}^{k,l}$ implicitly select the corresponding branches (cf. Eq. (18c)–(18d)). Alternative partitioning rules are also valid; for instance, we also implemented a rule based on the objective gradient components corresponding to complementarity variables $x_{1,i}$ and $x_{2,i}$ Kirches2022 :
$\begin{split}&\mathcal{D}_{1}^{*}(x^{k}\!+\!d^{k,l})=\{i\in{\mathcal{I}}_{00}(x^{k}\!+\!d^{k,l})\!\mid\!\nabla f_{1,i}(x^{k}\!+\!d^{k,l})\geq\nabla f_{2,i}(x^{k}+d^{k,l})\},\\ &\mathcal{D}_{2}^{*}(x^{k}+d^{k,l})={\mathcal{I}}_{00}(x^{k}+d^{k,l})\setminus\mathcal{D}_{1}^{*}(x^{k}+d^{k,l}),\\ &{\mathcal{I}}^{*}_{1}(x^{k}+d^{k,l})={\mathcal{I}}_{0+}(x^{k}+d^{k,l})\cup\mathcal{D}_{1}^{*}(x^{k}+d^{k,l})\\ &{\mathcal{I}}^{*}_{2}(x^{k}+d^{k,l})={\mathcal{I}}_{+0}(x^{k}+d^{k,l})\cup\mathcal{D}_{2}^{*}(x^{k}+d^{k,l}).\end{split}$ (24)
Since the BNLPs are in general nonconvex problems, starting from $x^{k}$ can give different objective improvements $f(x^{k})-f(x^{k,l})>0$ for different BNLPs. For convergence, it is not necessary to pick the best or a particular one, but only sufficient decrease must be guaranteed by the NLP solver, cf. Theorem 4.2. In our numerical experiments, we did not observe performance differences between using rule (23) and  (24).
If $\rho^{k,l}>\bar{\rho}$, the partition may not change between iterations, which is why line 10 checks whether a new BNLP needs to be solved. If no new partition is selected, the trust-region radius is reduced (line 17). Moreover, for larger $\rho^{k,l}$, if the current BNLP does not improve the objective (cf. Example 3.2), the trust-region radius is reduced as well (line 13). In numerical experiments, we observed that allowing larger initial trust-region radii often accelerates convergence, though they are not necessary for convergence.
4 Convergence analysis
This section provides the convergence proof of the proposed method. Before stating the main results, we discuss the main assumption and some auxiliary results.
4.1 Assumptions
Since the feasible set of any BNLP is a subset of the MPEC’s feasible set, an unbounded MPEC can be detected via an unbounded BNLP. Most NLP solvers readily detect this. Hence, in our convergence analysis we restrict attention to bounded problems.
Assumption 4.1
There exists a compact set $X$ such that for the feasible set $\Omega$ of the MPEC (1) it holds that $\Omega\subset X$.
Some results in the previous section required only the MPEC-ACQ. For the convergence results in this section, we require a stronger assumption.
Assumption 4.2
Then the MPEC-MFCQ holds at any feasible point of the MPEC (1).
MPEC-MFCQ implies MPEC-ACQ (Flegel2005a, , Theorem 3.1), and therefore the degenerate case of Example 3.1 is avoided. Other useful consequences are as follows.
Proposition 4.1
Suppose that Assumption 4.2 holds at $x^{*}\in\Omega$ for (1). Then:
  1. 1.
    the usual MFCQ is satisfied at $x^{*}$ by the corresponding ${\mathrm{BNLP}}({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))$ for any partition $({\mathcal{I}}_{1}(x^{*}),{\mathcal{I}}_{2}(x^{*}))\in\mathcal{P}(x^{*})$,
  2. 2.
    there exists a neighborhood $\mathcal{N}$ of $x^{*}$ and a constant $\bar{\tau}>0$ such that the usual MFCQ for Reg($\tau$) is satisfied at all $x\in\mathcal{N}\cap\Omega_{\mathrm{Reg}(\tau)}$ for all $\tau\leq\bar{\tau}$, where $\Omega_{\mathrm{Reg}(\tau)}$ is the feasible set of Reg($\tau$); and,
  3. 3.
    independently of Assumption 4.2, the tangent cones of the MPEC and the BNLPs satisfy
    $\mathcal{T}_{\Omega}(x^{*})=\bigcup_{({\mathcal{I}}_{1}^{*},{\mathcal{I}}_{2}^{*})\in\mathcal{P}^{*}}\mathcal{T}_{\Omega_{{\mathrm{BNLP}}({\mathcal{I}}_{1}^{*},{\mathcal{I}}_{2}^{*})}}(x^{*}).$ (25)
    Moreover, under Assumption 4.2, the left-hand side can be replaced by the MPEC-linearized feasible cone $\mathcal{F}_{\Omega}(x^{*})$, and the terms on the right-hand side by the usual linearized feasible cones of the BNLPs; cf. Eq. (40).
Proof. The first fact is (Flegel2005a, , Lemma 3.3), the second is (Hoheisel2013, , Theorem 3.2), and the third follows from (Flegel2005a, , Lemma 3.1) and (Flegel2005a, , Theorem 3.1). ∎
Many MPEC methods require solving a sequence of NLPs, e.g., Reg($\tau$) with a decreasing sequence $\{\tau^{k}\}$. In their convergence analysis, it is typically assumed that a constraint qualification holds and that a stationary point of such an NLP is available, without further specifying what assumptions on the NLP solver are required to compute such a stationary point Hoheisel2013 ; Kanzow2015 ; Ralph2004 ; Steffensen2010 . If a minimizer, and not only a stationary point, is required, then a suitable MPEC second-order sufficient condition may be assumed. Here we proceed similarly by making assumptions on the convergence of the NLP solver.
Assumption 4.3 (NLP solver)
Consider a nonlinear program
$\min_{x}f(x)\quad\textnormal{s.t.}\quad g(x)\geq 0,$ (26)
where $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $g:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{g}}$ are twice continuously differentiable functions. Starting from an initial guess $x^{0}\in{\mathbb{R}}^{n}$, an NLP solver returns either:
  1. 1.
    a stationary point of (26); moreover, if $x^{0}$ is a feasible but not stationary point of (26), then it returns a point $x^{*}$ such that for some $\alpha>0$, it holds that:
    $f(x^{0})-f(x^{*})\geq\alpha>0,$ (27)
  2. 2.
    a stationary point of the feasibility problem $\min_{x}\|g^{-}(x)\|$.
This assumption is justified as follows. The NLP (26) covers all NLPs considered in this paper, including BNLPs (8) and (Reg($\tau$)), for which MFCQ holds by Assumption 4.2 and Proposition 4.1. Under the MFCQ and additional assumptions (e.g., boundedness of the Hessian, second-order sufficient conditions), convergence of stabilized sequential quadratic programming Wright2002 ; Wright1998 and interior-point methods Vicente2002 to a local minimizer is guaranteed.
The outcomes of Assumption 4.3 are met in practice by most globalized NLP solver implementations, e.g., IPOPT Waechter2006 , filterSQP (Fletcher2002b, , Theorem 7), SNOPT Gill2005 , KNITRO Byrd2006 , to name a few.
Note that the globalization strategies of all listed NLP solvers enforce sufficient decrease conditions on every step to guarantee convergence to a stationary point. When starting from a feasible point, we assume that the step towards a stationary point satisfies a sufficient decrease conditions, which we summarize in Assumption 4.3 by inequality (27). Even when starting from a feasible point, the NLP solver may fail to produce a point satisfying the globalization strategy (i.e., sufficiently feasible with improved objective). In such cases, a feasibility restoration phase is typically triggered Waechter2006 , possibly returning a point with higher objective. We did not observe this in any of our numerical experiments and therefore do not treat it further. A possible remedy is to add the constraint $f(x)\leq f(x^{k})+\epsilon$, with $\epsilon$ small, to the current BNLP.
Lastly, we comment on the convergence of an LPEC solver. We solve the LPEC (15) via its MILP reformulation (18). MILP solvers solve LPs or certify their infeasibility, e.g., within a branch-and-bound algorithm, which always converges Wolsey1998 . Both the interior-point method and the dual-simplex method (with suitable anti-cycling rules) either solve an LP or certify its infeasibility; see, e.g., Nocedal2006 . Thus, in our setting, the LPECs can always be solved or certified infeasible. We implement this using either the commercial solver Gurobi Gurobi or the open-source solver HiGHS Huangfu2018 , both based on branch-and-bound methods.
4.2 Main convergence results
We start with a useful proposition. A similar result is available in Kirches2022 ; Leyffer2007 .
Proposition 4.2
Let $x\in\Omega$ be a feasible point for the MPEC (1), satisfying Assumption 4.2, and assume that $x$ is not B-stationary in the sense of Definition 2.1. Then there exist $\epsilon>0$, a step $s\in{\mathbb{R}}^{n}$ with $\|s\|_{\infty}=1$, and a partition $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ of ${\mathcal{I}}_{00}(x)$ such that:
$\nabla f(x)^{\top}s<-\epsilon,$ (28a)
$\nabla c_{i}(x)^{\top}s\geq\epsilon,$ $i\in{\mathcal{A}}(x),$ (28b)
$s_{1,i}=0,$ $i\in{\mathcal{I}}_{0+}(x),$ (28c)
$s_{2,i}=0,$ $i\in{\mathcal{I}}_{+0}(x),$ (28d)
$s_{1,i}=0,\ s_{2,i}\geq 0,$ $i\in{\mathcal{D}}_{1}(x),$ (28e)
$s_{1,i}\geq 0,\ s_{2,i}=0,$ $i\in{\mathcal{D}}_{2}(x).$ (28f)
Proof. If $x$ is not B-stationary, then there exists at least one BNLP for which $x$ is not stationary. Denote by $({\mathcal{D}}_{1}(x),{\mathcal{D}}_{2}(x))$ the partition of ${\mathcal{I}}_{00}(x)$ associated with this BNLP. By Proposition 4.1, the usual MFCQ holds for this BNLP. From this we draw two conclusions.
First, since $x$ is not B-stationary for this BNLP, there exists a $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ such that $\nabla f(x)^{\top}d<0$. Moreover, because MFCQ holds, the tangent cone $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ coincides with the usual linearized feasible cone of the BNLP, rendering (4) into an LP.
Second, considering the active sets for the BNLP at $x$, MFCQ implies the existence of a nonzero direction $s$ that points into the strict interior of the LP’s (4) feasible set, i.e., the following set is nonempty:
$S=\{\,$ $s\in{\mathbb{R}}^{n}\mid\nabla c_{i}(x)^{\top}s>0,\ i\in{\mathcal{A}}(x);$
$s_{1,i}=0,\ i\in{\mathcal{I}}_{0+}(x);\ s_{2,i}=0,\ i\in{\mathcal{I}}_{+0}(x);$
$s_{1,i}=0,\ s_{2,i}>0,\ i\in{\mathcal{D}}_{1}(x);\ s_{1,i}>0,\ s_{2,i}=0,\ i\in{\mathcal{D}}_{2}(x)\,\}.$
Next, let $\bar{s}\in S$. If $\nabla f(x)^{\top}\bar{s}<0$, set $s:=\bar{s}$. Otherwise, pick any $d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ with $\nabla f(x)^{\top}d<0$. Because $\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$ is a cone then for every $\lambda>0$ we can define a new descent direction $d_{\lambda}:=\lambda d\in\mathcal{T}_{\Omega_{\mathrm{BNLP}}}(x)$. Next, define the vector $s(t):=\bar{s}+td_{\lambda}$. Since $\bar{s}\in S$, there exists $t_{0}>0$ such that $s(t)\in S$ for all $t\in[0,t_{0}]$. Moreover,
$\nabla f(x)^{\top}s(t)=\nabla f(x)^{\top}\bar{s}+t\,\nabla f(x)^{\top}d_{\lambda}=\nabla f(x)^{\top}\bar{s}+t\lambda\,\nabla f(x)^{\top}d.$
Choose $\lambda$ so large that $0<\frac{\nabla f(x)^{\top}\bar{s}}{-\lambda\nabla f(x)^{\top}d}<t_{0}$, and then take any $t\in\big(\frac{\nabla f(x)^{\top}\bar{s}}{-\lambda\nabla f(x)^{\top}d},\,t_{0}\big]$. This yields $s(t)\in S$ and $\nabla f(x)^{\top}s(t)<0$. Fix such $(\lambda,t)$ and set $s:=s(t)$.
Because inequalities are strict for $s(t)$, define $\epsilon_{1}:=-\nabla f(x)^{\top}s,\ \epsilon_{2}:=\min_{i\in{\mathcal{A}}(x)}\nabla c_{i}(x)^{\top}s,\ \epsilon_{3}:=\min(\min_{i\in{\mathcal{D}}_{1}(x)}s_{2,i},\!\min_{i\in{\mathcal{D}}_{2}(x)}s_{1,i}),$ which are all positive, and set $\bar{\epsilon}:=\min\{\epsilon_{1},\epsilon_{2},\epsilon_{3}\}>0$. With $\hat{s}:=s/\|s\|_{\infty}$ and $\epsilon:=\bar{\epsilon}/\|s\|_{\infty}$, the vector $\hat{s}$ satisfies (28).
Next, we show that an LPEC constructed at an almost feasible point can “predict” a feasible BNLP. In our implementation, such a point, and the corresponding LPEC are found in Algorithm 2.
Theorem 4.1
Let Assumptions 4.1-4.3 hold. Consider $x^{*}\in\Omega$ a feasible point of the MPEC (1), and $\tilde{x}\in X$ (not necessarily an element of $\Omega$) such that $\|\tilde{x}-x^{*}\|\leq\delta$. Then, there exist positive constants $\kappa$, $\epsilon$ such that for a sufficiently small $\delta$ the following interval is nonempty
$\max\Big\{\frac{h_{c}(\tilde{x})}{\epsilon},h_{\perp}(\tilde{x})\Big\}\leq\rho\leq\kappa,$ (29)
and for any $\rho$ from this interval the $\mathrm{LPEC}(\tilde{x},\rho)$ in (15) is feasible. Moreover, for any feasible point $d$ of this LPEC, it follows that $\mathrm{BNLP}({\mathcal{I}}_{1}(\tilde{x}+d),{\mathcal{I}}_{2}(\tilde{x}+d))$ is feasible as well.
Proof. Consider the case when $\tilde{x}\in\Omega$, i.e., $\delta=0$ and $h_{c}(\tilde{x})=h_{\perp}(\tilde{x})=0$. The $\mathrm{LPEC}(\tilde{x},\rho)$ is always feasible for this point and let $d$ be an element of its feasible set. Now set $\kappa=\bar{\rho}$, where $\bar{\rho}$ is defined in (16). Then according to Lemma 3.2, for all $\rho\in(0,\kappa)$, every ${\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+d),{\mathcal{I}}_{2}(\tilde{x}+d))$ is feasible.
Next, we prove the assertion for $\tilde{x}\notin\Omega$. The main idea of the proof is to consider a feasible $x^{*}\in\Omega$, and to show that Lemma 3.2 is also true for points $\tilde{x}$ close enough to $x^{*}$. Denote by $\mathcal{B}(x,\rho)=\{d\in{\mathbb{R}}^{n}\mid\|d-x\|_{\infty}\leq\rho\}$ the $\ell_{\infty}$-ball centered at $x\in{\mathbb{R}}^{n}$ with the radius $\rho>0$.
First, we show that LPECs defined at points close to $x^{*}$ are also feasible under suitable conditions. Let $x^{*}\in\Omega$ be feasible but not B-stationary. It follows form Proposition 4.2 and the continuity of the functions $f$, and $c$, that there exists a neighborhood $\mathcal{N}^{*}=\mathcal{B}(x^{*},\kappa_{1})\cap X$, such that for all $\tilde{x}\in\mathcal{N}^{*}$ it holds that:
$\nabla f(\tilde{x})^{\top}d<-\epsilon,\nabla c_{i}(\tilde{x})^{\top}s\geq\epsilon,\ i\in{\mathcal{A}}(x^{*}).$ (30)
Next, we prove that every ${\mathrm{LPEC}}(\tilde{x},\rho)$ for all $\tilde{x}\in\mathcal{N}^{*}$ and a suitable $\rho$ is feasible. We start first with the general inequality constraints and use an argument similar to (Chin2003, , Lemma 4). Take the step $d=\rho s$, which satisfies $\|d\|\leq\rho$, and consider the linearization of the inequality constraints active at $x^{*}$:
$c_{i}(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq-h_{c}(\tilde{x})+\rho\epsilon,$ $i\in{\mathcal{A}}(x^{*}).$ (31)
The first term is lower bounded by the infeasibility of the inequality constraints, the second uses the fact that $\|s\|_{\infty}=1$ and (30). For the inequality constraints inactive at $x^{*}$, it follows from the continuity of $c$ that there exists positive constants $\gamma_{1}$ and $\gamma_{2}$ independent of $\tilde{x}$ such that $c_{i}(\tilde{x})\geq\gamma_{1}$ and $\nabla c_{i}(\tilde{x})^{\top}s\geq-\gamma_{2}$. Thus for the inactive constraints, we have
$c_{i}(\tilde{x})+\rho\nabla c_{i}(\tilde{x})^{\top}s\geq\gamma_{1}-\rho\gamma_{2},$ $i\notin{\mathcal{A}}(x^{*}).$ (32)
By requiring that the right-hand-side of (31) and (32) are nonnegative, we find that the linearized inequality constraints are feasible if $\rho$ satisfies:
$\frac{h_{c}(\tilde{x})}{\epsilon}\leq\rho\leq\frac{\gamma_{1}}{\gamma_{2}}.$ (33)
We now consider the complementarity constraints: $0\leq\tilde{x}_{1}+d_{1}\perp\tilde{x_{2}}+d_{2}\geq 0$. It can be seen that to make these constraints always feasible the trust region radius $\rho$ must be within the following bounds:
$\max(|\min(\tilde{x}_{1},\tilde{x}_{2}))|)\leq\rho\leq\max(\max(|\tilde{x}_{1}|,|\tilde{x}_{2}|)).$
[…]
$\delta\leq\rho\leq\kappa_{2}$ (34)
Therefore, by combining (33) and (34) we have that for all $\tilde{x}\in\mathcal{N}^{*}$ and $\rho$ for which holds:
$\max\Big\{\frac{\delta}{\epsilon},\delta\Big\}\leq\rho\leq\min\Big\{\kappa_{2},\frac{\gamma_{1}}{\gamma_{2}}\Big\}.$ (35)
Since there always exists a $\varepsilon>0$ for every $x^{*}\in\Omega$ (see proof of Proposition 4.2), the term $\lim\limits_{\delta\to 0}\frac{\varepsilon}{\delta}=0$ is well-defined. Therefore, the ${\mathrm{LPEC}}(\tilde{x},\rho)$ is feasible, since for a sufficiently small $\delta$ the interval (35) is nonempty.
In the final step of the proof, we make sure that:
$\mathcal{B}(\tilde{x},\rho)\subset\mathcal{B}(x^{*},\bar{\rho}).$ (36)
This can be achieved by computing:
$\|x-x^{*}\|_{\infty}\leq\|x-\tilde{x}\|_{\infty}+\|\tilde{x}-x^{*}\|_{\infty}\leq\delta+\rho\leq\bar{\rho},$
yielding the upper bound on $\rho\leq\bar{\rho}-\delta$. Combining this with (35) we require that
$\max\Big\{\frac{\delta}{\epsilon},\delta\Big\}\leq\rho\leq\min\Big\{\kappa_{2},\frac{\gamma_{1}}{\gamma_{2}},\bar{\rho}-\delta\Big\}.$ (37)
For sufficiently small $\delta$ this interval is nonempty and we define $\kappa=\min\Big\{\kappa_{2},\frac{\gamma_{1}}{\gamma_{2}},\bar{\rho}-\delta\Big\}$.
Furthermore, we conclude the following about the feasible ${\mathrm{LPEC}}(\tilde{x},\rho)$. Let $\tilde{d}$ be any feasible point of ${\mathrm{LPEC}}(\tilde{x},\rho)$. Due to (36), all points $\tilde{x}+\tilde{d}$ that can be reached from $\tilde{x}$ and that satisfy the complementarity constraints, can also be reached via $x^{*}+d$, where $d$ is a feasible point of ${\mathrm{LPEC}}(x^{*},\bar{\rho})$. Because Lemma 3.2 holds for ${\mathrm{LPEC}}(x^{*},\bar{\rho})$, we conclude that also all $){\mathrm{BNLP}}({\mathcal{I}}_{1}(\tilde{x}+\tilde{d}),{\mathcal{I}}_{2}(\tilde{x}+\tilde{d}$)) are feasible. Fig. 4 illustrates the arguments. This concludes the proof. ∎
(a) Illustration for some $i\in{\mathcal{I}}_{+0}(x^{*})$
(b) Illustration for some $j\in{\mathcal{I}}_{00}(x^{*})$
Figure 4: Illustration of the argument used the proof of Theorem 4.1.
Remark 4.1
Theorem 4.1 ensures that any algorithm that produces a point sufficiently close to the MPEC feasible set can be used in Phase I. Subsequently, finding a feasible point of the LPEC constructed at this point helps identify a feasible BNLP. In Section 3.3, we introduced a specific Phase I algorithm and, in the following corollary, we specialize Theorem 4.1 to this algorithm.
Corollary 4.1
Let Assumptions 4.1-4.3 hold. Let $x^{*}(\tau)$ be a stationary point of $\mathrm{Reg}(\tau)$ for some $\tau>0$. There exist positive constants $\bar{\tau}$, $\kappa$ such that for any: $\tau\in\left(0,\bar{\tau}\right]$ and $\rho\in[\sqrt{\tau},\kappa]$ the $\mathrm{LPEC}(x^{*}(\tau),\rho)$ in (15) is feasible. Moreover, for any feasible point $d$ of this LPEC, it follows that $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$ is feasible as well.
Proof. First, regard the case when $x^{*}(\tau)\in\Omega$, for some $\bar{\tau}\geq\tau>0$. The $\mathrm{LPEC}(x^{*}(\tau),\rho)$ is always feasible for this point and let $d$ be an element of its feasible set. Now for $\kappa=\bar{\rho}$ and all $\rho\in(0,\kappa)$, according to Lemma 3.2 all $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$ are feasible.
Second, for $x^{*}(\tau)\notin\Omega$ set $\tilde{x}=x^{*}(\bar{\tau})$ and apply Theorem 4.1. Since $x^{*}(\bar{\tau})$ is a stationary point of $\mathrm{Reg}(\bar{\tau})$, it holds that $c(x^{*}(\bar{\tau}))\geq 0$, i.e., $h_{c}(\tilde{x})=0$. Moreover, from $x_{1,i}x_{2,i}\leq\tau,\ \forall i\in\{1,\ldots,m\}$, it follows that $h(\tilde{x})\leq\sqrt{\bar{\tau}}=\delta$. Hence, for a sufficiently small $\delta=\bar{\tau}$ and every $0<\tau\leq\bar{\tau}$ the interval in (37) is nonempty and we can apply Theorem 4.1. ∎
Remarkably, it follows from the last two results, that the LPEC in Phase I does not even need to be solved to optimality. It is sufficient to find a feasible point of the LPEC to identify a feasible BNLP. Before we present the main convergence result, we prove that the inner loop in Algorithm 3 terminates in a finite number of iterations.
Proposition 4.3
[…]
Proof. If $d^{k,l}=0$ finite termination is trivial, so consider $d^{k,l}\neq 0$. Each step rejection in lines 10 or 12 multiplies $\rho$ by $\gamma^{\mathrm{L}}\in(0,1)$, hence after finitely many iterations there is an index $N\geq l$ with $\rho^{k,N}=(\gamma^{\mathrm{L}})^{N-l}\rho^{k,l}<\bar{\rho}$. For $\rho^{k,N}<\bar{\rho}$, due to Lemma 3.1 we can focus on the reduced LPEC (12). In this case, because $\|d^{k,N}\|_{\infty}<\bar{\rho}$ is sufficiently small, from the proof of Lemma 3.2 we have
$\mathcal{P}(x^{k}+d^{k,N})\subseteq\mathcal{P}(x^{k}),$
so every BNLP that can be selected at $x^{k}+d^{k,N}$ can also be selected at $x^{k}$, hence $x^{k}$ is feasible for all these BNLPs.
Since no termination occurred before index $N$, it holds that $\nabla f(x^{k})^{\top}d^{k,N}<0$ and $d^{k,N}\neq 0$. By (10) the vector $d^{k,N}$ solves all branch LPs of the reduced LPEC, and by (40) each such branch LP corresponds to an LP of the form (4) for the BNLPs. Therefore $d^{k,N}\neq 0$ is a common minimizer of all these LPs with $\nabla f(x^{k})^{\top}d^{k,N}<0$, which implies that $x^{k}$ is feasible but not stationary for any BNLP indexed by $({\mathcal{I}}_{1}(x^{k}+d^{k,N}),{\mathcal{I}}_{2}(x^{k}+d^{k,N}))\in\mathcal{P}(x^{k}+d^{k,N})$.
By Assumption 4.3, the NLP solver applied to any such BNLP produces a point $x^{k,N}$ with $f(x^{k,N})<f(x^{k})$. Hence an acceptance occurs and the inner loop terminates. Since $N$ is finite, the inner loop terminates after finitely many iterations. ∎An immediate consequence of this proposition is that we can terminate an LPEC solver early at a nonzero feasible point $d\neq 0$.
Corollary 4.2
[…]
Proof. It follows from the proof of the previous proposition that, for sufficiently large $N$, the solution of the LPEC, $d^{k,N}\neq 0$, also solves every branch LP. By item (3) of Proposition 4.1, and for $\rho^{k,N}<\bar{\rho}$, the feasible set of a branch LP is a convex cone intersected with the trust-region constraints. Because it is a cone, from the feasibility of $\hat{d}=0$ and $d^{k,N}$ it follows that any $d=td^{k,N}$ with $t>0$ is also feasible (within the trust region), and $\nabla f(x^{k})^{\top}(td^{k,N})=\nabla f(x^{k})^{\top}d<0$. Moreover, because $d=t\ d^{k,N}$ it holds that $(\mathcal{I}_{1}(x^{k}+d),\mathcal{I}_{2}(x^{k}+d))$ coincides with $(\mathcal{I}_{1}(x^{k}+d^{k,N}),\mathcal{I}_{2}(x^{k}+d^{k,N}))$. ∎
Theorem 4.2
Regard the MPEC (1) and suppose that Assumptions 4.1-4.3 hold. Then MPECopt given in Algorithm 3, with a Phase I from Section 3.3 has one of the following mutually exclusive outcomes:
  1. 1.
    In Phase I, a stationary point of the problem of minimizing a measure of constraint infeasibility is returned.
  2. 2.
    It computes a point $x^{*}$ that is B-stationary for the MPEC (1) after solving finitely many BNLPs and LPECs.
  3. 3.
    Create an infinite sequence of iterates $\{x^{k}\}$ with decreasing objective values, whose accumulation point is B-stationary.
If, in addition, every stationary point $\hat{x}$ of each BNLP is isolated, i.e., there exists a neighborhood $\hat{\mathcal{N}}$ of $\hat{x}$ containing no other stationary points, then outcome 3 cannot occur.
[…]
Given a feasible point $x^{0}$, we discuss the outcomes of Algorithm 3. If follows from Proposition 4.3 that the inner loop terminates in a finite number of iterations. Therefore, the outer loop it produces a sequence of points $\{x^{k}\}\subset\Omega$ with decreasing objective values $\{f^{k}\}$. Consider some iterate $x^{k}$, which is a stationary point of some $\mathrm{BNLP}({\mathcal{I}}_{1}(x^{k}),{\mathcal{I}}_{2}(x^{k}))$, and a $\rho^{k,l}>0$. If $d^{k,l}=0$ is a minimizer of $\mathrm{LPEC}(x^{k},\rho^{k,l})$, then a B-stationary point is found and the algorithm terminates in a finite number of iterations, which is outcome two.
We prove the occurrence of outcome three. Since $\{x^{k}\}\subset\Omega\subset X$ and $f$ is continuous while $X$ is compact, the sequence $\{f(x^{k})\}$ is bounded. By monotonicity, $\{f(x^{k})\}$ is convergent to some $f^{\infty}\in{\mathbb{R}}$. Consequently, $\{x^{k}\}$ has an accumulation point $x^{\infty}\in X$. Let $\{x^{k_{j}}\}$ be a subsequence converging to $x^{\infty}$. Suppose, for contradiction, that $x^{\infty}$ is not a B-stationary point. Recall that each $x^{k},k>0$ is a stationary point of a BNLP, and by Assumption 4.3, for each $k\geq 0$ there exist an $\alpha^{k}>0$ such that $f(x^{k})-f(x^{k+1})\geq\alpha^{k}>0.$ Therefore, we can define an uniform bound $\bar{\alpha}:=\min_{k}\alpha^{k}>0$. It follows that for the subsequence, $f(x^{k_{j}})-f(x^{k_{j+1}})\;\geq\;\bar{\alpha},\quad\forall j\in\mathbb{N}.$ Summing over $j$ yields the telescoping bound
$f(x^{k_{j}})=f(x^{k_{0}})-\sum_{\ell=0}^{j-1}\big(f(x^{k_{\ell}})-f(x^{k_{\ell+1}})\big)\;\leq\;f(x^{k_{0}})-j\bar{\alpha}.$
Letting $j\to\infty$ gives $f(x^{k_{j}})\to$-$\infty$. However, since $x^{k_{j}}\to x^{\infty}\in X$ and $f$ is continuous on the compact set $X$, the values $\{f(x^{k_{j}})\}$ must remain bounded. This contradiction our assumption, and therefore $x^{\infty}$ is a B-stationary point.
[…]
We note that the assumption of isolated stationary points for finite termination is not overly restrictive. Under MPEC-MFCQ in Ass. 4.2, the usual MFCQ holds for the associated BNLPs. Moreover, if we additionally assume generalized second-order sufficient conditions at the BNLP stationary points, these points are isolated local minimizers (Forsgren2002, , Theorem 6.9).
5 Numerical results
In this section, we extensively compare several MPECopt variants with regularization-based methods, as well as with solving the MINLP and NLP reformulations of MPECs on two distinct nonlinear MPEC benchmarks.
5.1 Implementation
In our implementation and experiments, we treat MPECs more general than (1), given in the form of:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $f(x)$ (38a)
s.t. $c^{\mathrm{lb}}\leq c(x)\leq c^{\mathrm{ub}},$ (38b)
$x^{\mathrm{lb}}\leq x\leq x^{\mathrm{ub}},$ (38c)
$0\leq g(x)\perp h(x)\geq 0,$ (38d)
where $c^{\mathrm{lb}},c^{\mathrm{ub}}\in{\mathbb{R}}^{n_{c}}$ are lower and upper bounds of $c(x)$ with $c^{\mathrm{ub}}_{i}\geq c^{\mathrm{lb}}_{i},\ \forall i\in\{1,\ldots,n_{c}\}$. Equality constraints are imposed by settings $c^{\mathrm{ub}}_{i}=c^{\mathrm{lb}}_{i}$. The vectors $x^{\mathrm{lb}},x^{\mathrm{ub}}\in{\mathbb{R}}^{n}$ are lower and upper bounds on $x$. If there is no lower or upper bound on some $x_{i}$ or $c_{i}(x)$, then we set the respective value on $\texttt{inf}$ or inf. The functions $g,h:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ are assumed twice continuously differentiable.
To solve the NLPs arising in MPECopt we use the IPOPT solver Waechter2006 . For better performance, we change some of the default options by setting the tolerance to $\texttt{tol}=10^{-12}$ (default is $10^{-8}$), $\texttt{acceptable\_tol}=10^{-9}$ (default is $10^{-6}$), and $\texttt{warm\_start\_init\_point}=\texttt{'yes'}$. Moreover, we change the barrier parameter strategy by setting: mu_strategy = ’adaptive’ and mu_oracle = ’quality-function’, cf. Nocedal2009 . In all experiments, we use MA27 Duff1982 from the HSL library HSL as the linear solver in IPOPT. The same settings are used for all methods and NLP solutions in the following experiments.
All problems are formulated in the symbolic framework of CasADi Andersson2019 via its Matlab interface. CasADi provides the first and second-order derivatives for all problem functions via its automatic differentiation and has an interface for calling IPOPT.
The LPECs are solved via the MILP formulation in (18) and solve it with the commercial solver Gurobi Gurobi or the open-source solver HiGHS Huangfu2018 . In our experiments, the MILP reformulation was significantly faster that regularization- and penalty-based methods, cf. Appendix C. For the MILPs, we set the maximum number of nodes in the branch and bound algorithm to 5000 and the maximum computation time to 5 minutes. The tolerance for the LP solves in both MILP solvers is set to $10^{-9}$ (default is $10^{-6}$).
For comparisons, we also implement the regularization and penalty-based methods. Thereby, an adaptation of Reg($\tau$), $\ell_{1}$ and $\ell_{\infty}$ Kim2020 ; Ralph2004 ; Nurkanovic2024b to the form of  (38) are solved in a homotopy loop with a decreasing parameter $\tau$. We initialize $\tau^{0}=1$ and update it using the rule $\tau^{k}=0.1\tau^{k-1}$. The loop terminates if either: the NLP solution is feasible and the complementarity tolerance $h_{\perp}(x)$ is below $10^{-9}$; NLP is locally infeasible; or a maximum number of $15$ homotopy iterations is reached. A homotopy loop is declared successful if it finds a stationary point. Additionally, we compare MPECopt to solving the MPEC directly as an NLP in (2) with IPOPT, and solving its MINLP reformulation (21) with a nonlinear branch-and-bound method implemented in Bonmin Bonami2008 .
We mention the remaining parameter choices in our implementation. The parameters in Algorithm 2 are $\kappa=0.1$, $\tau^{0}=1$, ${\rho}^{I}=10^{-1}$, and $\texttt{tol}_{h}=10^{-10}$. In the implementation of Algorithm 3, in line 6, a point is declared to be B-stationary if $\|d^{k,l}\|_{\infty}\leq\texttt{tol}_{\mathrm{B}}$. However, since there is also the trust region constraint $\|d\|\leq\rho^{k,l}$ we should avoid having $\rho^{k,l}<\texttt{tol}_{\mathrm{B}}$. Since the inner loop terminates within a finite number of iterations (cf. Proposition 4.3), this situation can be avoided by choosing appropriate parameters of the algorithm; we pick $\texttt{tol}_{\mathrm{B}}=10^{-8}$, $\rho^{k,0}=10^{-3}$, $N^{\mathrm{in}}=6$ and $\gamma^{\mathrm{L}}=0.1$. We also set a maximum number of outer iterations to $N^{\mathrm{out}}=25$. If no B-stationary point is found, the algorithm terminates after reaching the maximum number of iterations.
Our implementation of MPECopt is open-source and available at github.com/nosnoc/mpecopt. It contains code the reproduce the two benchmark results below, and several tutorial examples.
5.2 Results on MacMPEC benchmark
The MacMPEC problem set is available in the form of a package of AMPL .mod and .dat files from wiki.mcs.anl.gov/leyffer/index.php/MacMPEC. We consider in total 191 problems. Results of the experiments are given in the Dolan-Moré absolute and relative performance profiles Dolan2002 .
(a) Absolute timings.
(b) Relative timings.
Figure 5: Evaluating different MPEC solution methods on the MacMPEC test set in terms of finding a B-stationary point.
We compare MPECopt to other mentioned approaches. For Phase I of MPECopt, we use Algorithm 2 with Gurobi as LPEC solvers, as it had the highest success compared to other variants we tested. In the experiments, we compare in total six algorithms:
  1. 1.
    MPECopt, with regularization-based Phase I, and Gurobi as LPEC solver – MPECopt-Reg-Gurobi.
  2. 2.
    MPECopt, with $\ell_{\infty}$-penalty-based Phase I, and Gurobi as LPEC solver – MPECopt-$\ell_{\infty}$-Gurobi.
  3. 3.
    Regularization method that solve Reg($\tau$) in a homotopy loop – Reg.
  4. 4.
    $\ell_{\infty}$-penalty method, that solve an $\ell_{\infty}$ reformulation of (1Kim2020 in a homotopy loop – $\ell_{\infty}$.
  5. 5.
    Directly solving the MPEC as an NLP (2) with IPOPTNLP.
  6. 6.
    Solving the MINLP reformulation (21) with a nonlinear branch and bound in BonminMINLP.
We also tried the $\ell_{1}$ penalty method, which has a similar computation time as $\ell_{\infty}$, but a slightly lower success rate. Moreover using HiGHS instead of Gurobi lead to slightly slower method and marginally lower success rate. Therefore, for better readability, we omit these variants from the plots. Since all methods use IPOPT as the NLP solver, we can directly compare the computation times. Fig. 5 reports the absolute and relative performance profiles for this experiment.
Note that, if MPECopt is successful, it always returns a B-stationary point, whereas for the other methods, we can only know that it is a B-stationary point if it is also S-stationary. To verify B-stationarity of the solutions computed by all other methods, we proceed as described in Appendix B. We do not add the computational cost of verifying B-stationarity to the overall computation time of the other methods.
We observe that MPECopt-Reg-Gurobi achieves the highest overall success rate of 94.24%. It is both faster and more successful than its counterpart Reg (93.2%). The $\ell_{\infty}$-penalty method is the fastest overall and faster than MPECopt-Reg-Gurobi, but it attains a lower success rate of 92.14%. Meanwhile, MPECopt-$\ell_{\infty}$-Gurobi is only slightly slower than $\ell_{\infty}$ and achieves the same success rate. As expected, the plain NLP and MINLP reformulations are less competitive than the MPEC-tailored methods, with success rates of 90.00% and 72.52%, respectively. The main reasons for failure are timeouts and unsuccessful restoration phases. Additionally, the MINLP method frequently terminates at feasible but non–B-stationary points (39 cases).
The MPECopt method inherits the efficacy of the standard methods used in Phase I but can often be even faster, as it requires solving fewer NLPs. This is because, already for moderate values of $\tau$, the LPEC can identify a feasible BNLP in Phase I. Consequently, MPECopt needs to solve fewer NLPs than Reg and $\ell_{\infty}$ in most cases. The branch-and-bound MINLP method solves the largest number of NLPs, which also explains its slow computation times.
In practice, only a few LPECs need to be solved. For instance, on the MacMPEC benchmark, the number of LPECs solved ranged from 1 to 6. In many cases, only two are required: one in Phase I to identify a feasible BNLP, and one in Phase II to verify B-stationarity.
We have shown in Section 4 that, unless we need to verify B-stationarity, the LPECs do not need to be solved to global optimality; finding a feasible point is sufficient. Moreover, Proposition 3.1 shows that if the solution is S-stationary, it can be verified by solving a single LP. We examine these theoretical results by plotting histograms of the number of nodes in a MILP call (corresponding to the number of LPs solved) in both Phase I and Phase II. Fig. 6a shows the histogram when the MILP is solved to global optimality, while Fig. 6b shows the case when it is terminated early, at a feasible point. In both cases, Phase II only needed to solve LPs, and in Phase I we observe a significant reduction in the number of LPs to be solved in several cases. In fact, with early termination, on MacMPEC we only need to solve LPs. This illustrates our theoretical results: early termination of the MILP method reduces the number of LPs that must be solved. This did not significantly affect the overall computation time, as the largest computational burden on the MacMPEC problem set came from solving the NLP subproblems.
(a) Solving to optimality.
(b) Early termination.
Figure 6: Number of nodes in the MILP solver in Phase I and $I$. The node counter on the horizontal axis is on logarithmic scale.
In summary, all methods – sometimes with minor tuning – can successfully compute B-stationary points on most MacMPEC problems, although only MPECopt directly certifies this. All methods are quite fast on easy problems without biactive constraints, where often a single NLP solve is sufficient. In fact, more than two-thirds of the solutions found by any algorithm correspond to problems with an empty biactive set. To provide a broader perspective, we present another benchmark with larger problems in the next section.
(a) Absolute timings.
(b) Relative timings.
Figure 7: Evaluating different MPEC solution methods on the synthetic test set in terms of finding a stationary point.
5.3 Synthetic nonlinear MPEC benchmark
Many MacMPEC problems are small or easily solved by all examined algorithms. To further compare MPECopt, we use a synthetic nonlinear MPEC benchmark described in Appendix D. The benchmark consists of 75 significantly larger problems, ranging from 200 complementarity constraints and 500 variables in the smallest instances to 4000 complementarity constraints and 10000 variables in the largest. These problems are specifically designed to violate the MPEC-LICQ. In this experiment, we exclude the MINLP from the comparison, as it was already not competitive on the easier problem set. Fig. 7 shows the relative and absolute performance profiles of the methods we compared.
(a) Solving to optimality.
(b) Early termination.
Figure 8: Number of nodes in the MILP solver in Phase I and $I$. The node counter on the horizontal axis is on logarithmic scale.
Similar to the results on MacMPEC (Fig. 5), we see that the MPECopt variants are the fastest and most successful algorithms even on the large problem instances. For the same reasons as in the previous experiments, they require fewer total NLP solves. Moreover, on this set, the differences in computation times between MPECOpt-Reg-Gurobi (96.16% success rate), MPECOpt-$\ell_{\infty}$-Gurobi (94.67% success rate), and Pen-$\ell_{\infty}$ (94.67% success rate) are more pronounced, whereas Reg (94.67% success rate) needs to solve more NLPs and is therefore slower. The lower success rates of the regularization- and penalty-based methods compared to their MPECopt counterparts are due to either reaching the maximum number of iterations or hitting the time limit. In other cases, the methods failed by converging to a point of local infeasibility, i.e., when the restoration phase failed. Directly solving the NLP is by far the least successful approach (16.00% success rate), and in most cases it timed out after reaching the 45-minute limit. This further justifies the development of tailored methods for solving MPECs. As shown in Fig. 7a, the MPEC-tailored methods compute solutions considerably faster and rarely run into timeouts.
Lastly, we examine the effects of early termination in the LPEC solvers in Fig. 8. In Fig. 8a, where the MILP (and thus the LPEC) is solved to global optimality, we see that in both phases there are instances requiring a significant number of LPs. By contrast, Fig. 8b shows the effects of early termination in Phase I (due to Corollary 4.1) and Phase II (due to Corollary 4.2), where the number of LPs is drastically reduced. In particular, in Phase I a single LP was always sufficient. In Phase II, we also observe a reduction. We terminate the MILP solver if a feasible point with an objective below $0.1\|\nabla f(x^{k})\|\rho^{k,l}$ is found. With more careful tuning of this factor, the number of LPs in Phase II can often be reduced further.
In conclusion, we can see that the MPECopt is competitive with other existing approaches. Unlike all other methods, it always provides a certificate of B-stationarity. Moreover, the experiments demonstrate that solving an LPEC was never a computational bottleneck, even for the larger problem instances. Significant savings can be achieved through early termination, in line with our theoretical predictions.
6 Conclusions and summary
This paper introduces MPECopt, a two-phase method for computing B-stationary points of mathematical programs with complementarity constraints (MPECs). The first phase identifies a feasible point of the MPEC or certifies local infeasibility. The second phase iterates over feasible points to find a B-stationary point by solving a finite sequence of branch nonlinear programs (BNLPs) and linear programs with complementarity constraints (LPECs). We perform a detailed convergence analysis of the algorithm.
We also provide an open-source implementation of MPECopt, available at github.com/nosnoc/mpecopt. The implementation leverages robust and mature NLP and MILP solvers. We tested MPECopt on two numerical benchmarks: the MacMPEC set and a new synthetic nonlinear MPEC test set containing problems with up to four thousand complementarity constraints. The numerical results confirm the efficacy of the proposed method for finding B-stationary points. MPECopt is generally faster and more reliable than existing methods, requiring fewer NLP solves than homotopy-based approaches.
In general, LPECs are challenging combinatorial optimization problems that can be formulated as mixed-integer linear programs (MILPs). Interestingly, a combinatorial number of LPs only needs to be solved if the limit point is B-stationary but not S-stationary. We prove that in all other cases, the LPEC solver can be terminated early or that solving a single LP is sufficient. In our numerical experiments on the two benchmarks, even for large-scale problem instances, it was often sufficient to solve only the root node of the MILP reformulation of the LPEC, which corresponds to a single LP solve.
An intriguing observation from the benchmarks is that the limit point of regularization- or penalty-based methods is in most cases B-stationary, even when it is not S-stationary. To our knowledge, current convergence results for these methods do not explain this phenomenon. However, verifying B-stationarity typically requires solving additional nonlinear programs and LPECs, whereas MPECopt inherently certifies it, since solving LPECs is fully integrated into the algorithm.
Finally, in this paper we assume that the BNLPs are solved to full convergence. A natural direction for future work is to study inexact BNLP solutions while maintaining robust and fast convergence.
Acknowledgments
A.N. acknowledges the generous support of this research by the DFG via Research Unit FOR 2401, project 424107692, 504452366 (SPP 2364), and 525018088, by BMWK via 03EI4057A and 03EN3054B, and by the EU via ELO-X 953348. Moreover, A.N. thanks Anton Pozharskiy for translating MacMPEC to CasADi, and for many useful remarks on the MPECopt and benchmark implementations. This work was also supported by the Applied Mathematics activity within the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under contract number DE-AC02-06CH11357.
Data availability
All data necessary to reproduce the results in this paper are publicly available in the GitHub repository https://github.com/nosnoc/mpecopt. This repository contains the code, experiment scripts, and data used to generate the numerical results reported in the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Proofs of lemmas from Section 3
Proof of Lemma 3.1: Denote by $\Omega^{\mathrm{R}}$ and $\Omega^{\mathrm{F}}$ the feasible sets of the reduced LPEC (12) and full LPEC (15), excluding the constraints $\|d\|\leq\rho$, respectively. Compared to $\Omega^{\mathrm{R}}$, the set $\Omega^{\mathrm{F}}$ is additionally restricted by the linear inequalities for all inactive constraints, i.e., $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,i\in\bar{{\mathcal{A}}}^{k}$. Recall that the constraints $0\leq x_{1,i}^{k}+d_{1,i}\perp x_{1,i}^{k}+d_{2,i}\geq 0$ define the set $\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\cup\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$. Moreover, $\Omega^{\mathrm{F}}$ contains all branches excluded for the strictly active complementarity indices $i\in{\mathcal{I}}_{+0}^{k}\cup{\mathcal{I}}_{0+}^{k}$ in $\Omega^{\mathrm{R}}$. Therefore, the two sets are related as follows:
$\begin{split}\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}&\bigcap_{i\in\bar{{\mathcal{A}}}^{k}}\{d\mid c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0\}\\ &\bigcup_{i\in{\mathcal{I}}_{0+}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\\ &\bigcup_{i\in{\mathcal{I}}_{+0}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}.\end{split}$ (39)
To ensure that the two LPECs have the same set of solutions, we need to restrict $\|d\|_{\infty}\leq\rho$ such that $\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}$.
Let us first consider the inequality constraints. For all $i\in\bar{{\mathcal{A}}}^{k}$, we have $c_{i}(x^{k})>0$, and we want $c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d>0.$ If $\nabla c_{i}(x^{k})=0$, this inequality holds trivially for any $d$ (and thus any $\rho$). Otherwise, if $\nabla c_{i}(x^{k})\neq 0$, we can apply the Cauchy–Schwarz inequality $|\nabla c_{i}(x^{k})^{\top}d|\leq\|\nabla c_{i}(x^{k})\|\,\|d\|.$ Hence, the linearized inequality remains inactive as long as $\|d\|<\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}.$ To ensure this holds for all relevant $i$, we define $\bar{\rho}_{1}=\min_{i\in\bar{{\mathcal{A}}}^{k},\,\nabla c_{i}(x^{k})\neq 0}\frac{c_{i}(x^{k})}{\|\nabla c_{i}(x^{k})\|}.$ Next, we compute upper bounds for $\rho$ to exclude the union terms in (39).
Now observe that, for example, for any $i\in{\mathcal{I}}_{+0}^{k}$ it holds that $x_{1,i}^{k}>0$ and $x_{2,i}=0$. If $\rho$ is chosen such that $d_{1,i}<\rho<x_{1,i}^{k}$ then $x_{1,i}^{k}+d_{1,i}>0$ and $0\leq x_{1,i}^{k}+d_{1,i}\perp d_{2,i}\geq 0$ can be only satisfied for $d_{2,{i}}=0$. Thus, the branch $\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$ cannot be reached, i.e., it is infeasible. Similar reasoning holds for any $i\in{\mathcal{I}}_{0,+}^{k}$. Therefore, by picking $\rho<\bar{\rho}_{2}=\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\}$, none of the terms in the unions in (39) is feasible. The equation (16) follows from writing $\bar{\rho}=\min(\bar{\rho}_{1},\bar{\rho}_{2})$ explicitly. Thus, for $\rho<\bar{\rho}$, we have $\Omega^{\mathrm{F}}\cap\{\|d\|_{\infty}\leq\rho\}=\Omega^{R}\cap\{\|d\|_{\infty}\leq\rho\}$. As the two LPECs have the same objective, their sets of local minimizers coincide. ∎
Proof of Lemma 3.2: It follows from Lemma 3.1 that for $\rho\in(0,\bar{\rho})$ both the full and reduced LPEC have the same feasible set. Thus, we focus solely on the reduced LPEC. Denote by $\mathcal{P}^{k}$ the set of all possible partitions $(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})$ at $x^{k}\in\Omega$ (see Eq. (7) for definitions). The MPEC-MFCQ implies the MPEC-ACQ, hence the MPEC-linearized feasible cone and the standard linearized feasible cones of the BNLPs satisfy the following (Flegel2005a, , Lemma 3.1):
$\begin{split}\mathcal{F}_{\Omega}(x^{k})=\bigcup_{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})\in\mathcal{P}^{k}}\{d\in{\mathbb{R}}^{n}\mid&c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0,\ \forall i\in{\mathcal{A}}^{k},\\ &x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0,\ \forall i\in{\mathcal{I}}_{1}^{k},\\ &x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0,\ \forall i\in{\mathcal{I}}_{2}^{k}\}.\end{split}$ (40)
[…]
Appendix B Identifying stationary points
To determine the type of stationary point from Definition 2.3 we must solve the TNLP (9), which requires a correct identification of the active sets ${\mathcal{I}}_{+0},{\mathcal{I}}_{0+}$ and ${\mathcal{I}}_{00}$. MPECopt is an active set method where the complementarity constraints are usually satisfied exactly or close to machine precision, making it easy to find the active sets for the TNLP.
On the other hand, Reg, and Pen-$\ell_{1}$ rarely have a complementarity residual $h_{\perp}(x)$ much smaller than the prescribed tolerance of $10^{-9}$, making an active set identification more difficult. We take an iterative procedure, and identify the active set via:
${\mathcal{I}}_{0+}^{\epsilon}(x)=\{i\mid x_{1,i}\leq\epsilon,x_{2,i}>\epsilon\},$ (41a)
${\mathcal{I}}_{+0}^{\epsilon}(x)=\{i\mid x_{1,i}>\epsilon,x_{2,i}\leq\epsilon\},$ (41b)
${\mathcal{I}}_{00}^{\epsilon}(x)=\{i\mid x_{1,i}\leq\epsilon,x_{2,i}\leq\epsilon\},$ (41c)
with $\epsilon=10^{-6}$. To minimize the possibility of wrong conclusions, we implement several safeguards. We denote the solution computed via Reg or Pen-$\ell_{1}$ by $x^{*}$, and the corresponding TNLP solution by $x^{\mathrm{tnlp}}$. We declare the active set to be correctly identified if $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$. If this is not satisfied we reduce $\epsilon$ by a factor of $10$ and repeat the procedure. Once $\|x^{*}-x^{\mathrm{tnlp}}\|\leq 10^{-7}$ is true or a maximum number of 5 attempts is exceeded, we compute the multipliers in the TNLP and classify the stationary point. To verify B-stationarity, we take several steps. First, we use the point $x^{\mathrm{tnlp}}$ and construct an ${\mathrm{LPEC}}(x^{\mathrm{tnlp}},\rho)$ with $\rho=10^{-3}$. If the objective of this LPEC is below $\texttt{tol}_{\mathrm{B}}=10^{-8}$, then we declare a $x^{*}$ to be B-stationary. If this is not the case, we take up to three more LPEC solve attempts while reducing $\rho$ by a factor of 10. The minimal value is $\rho=10^{-6}$, which is the minimal allowed trust region radius in MPECopt. This delivers a valid certificate in most cases. To have more robustness to wrong active set identification, if this does not verify B-stationarity, we repeat the same but now with $x^{*}$ as linearization point ${\mathrm{LPEC}}(x^{*},\rho)$. Finally, if this also fails, we look at the objective $f(x^{*})$ and the objective computed by MPECopt for the same problem, which is denoted by f(x^)f($\hat${x}). If the relative difference satisfies $\frac{|f(x^{*})-f(\hat{x})|}{|f(\hat{x})|+10^{-16}}<10^{-3}$, then we declare $x^{*}$ to be B-stationary, if $\hat{x}$ is B-stationary. If the difference between the objective is larger, and the LPECs find a descent direction, the point $x^{*}$ is declared to be not B-stationary.
Appendix C Comparison of different LPEC solvers
We compare the cumulative NLP and LPEC computation times in every MPECopt solver call on every MacMPEC problem. In all experiments, we use the regularization-based Phase I described in Algorithm 2, but with different LPEC solvers in both phases. We compare solving the LPEC as MILP (18) and via regularization-based methods. The HiGHS solver performs significantly better if the trust region in the LPEC is smaller, thus we set $\rho_{I}=10^{-3}$ in Phase I, while Gurobi is less sensitive to $\rho_{I}$ values. We consider four variants: Gurobi-MILP ($\rho_{I}=10^{-1}$), HiGHS-MILP ($\rho_{I}=10^{-3}$), Reg-MPEC, and $\ell_{\infty}$-MPEC.
We solve all MacMPEC problems with four different LPEC solution strategies, comparing total cumulative computation times for solving all LPECs in each instance. In 34 of 191 problems, the MPEC is already solved or locally infeasible after the first NLP, so we exclude these (effective LPEC time is zero). The remaining 157 problems are solved by all approaches except HiGHS-MILP, which fails on siouxfls and siouxfls1 due to unsolved LPECs.
Fig. 9 shows the results, where the diagonal represents equal NLP and LPEC solving times. For MILP methods, most points fall below the diagonal, indicating that BNLPs dominate computation time rather than LPECs. For regularization methods, points cluster around/above the diagonal, showing computational burden is split almost equally between NLPs and LPECs. Specifically, cumulative NLP time exceeds LPEC time in 0% of cases for Gurobi-MILP, 20% for HiGHS-MILP, 87.26% for Reg-MPEC, and 85.99% for $\ell_{\infty}$-MPEC. MILP solvers benefit from smaller initial trust region radii in Phase I, but smaller $\rho_{I}$ requires solving more Phase I NLPs to find feasible LPECs, shifting computational load toward NLPs.
Figure 9: Comparison of total NLP and LPEC computation times on the MacMPEC.
Somewhat surprisingly, the global MILP algorithms significantly outperform the local regularization and penalty-based methods. The MILP-based methods are up to two orders of magnitude faster than regularization-based methods for LPECs on the MacMPEC test set. This difference can be explained by several observations. The MILP methods solve the LPEC to global optimality (sufficient for verifying B-stationarity), whereas the others compute only a stationary point (which may not be a local minimum or B-stationary point). Consequently, LPECs solved by regularization-based methods are more often locally infeasible in Phase I or require solving more LPECs until a B-stationary point is verified in Phase II. This confirms the arguments in Section 3.2.2 on efficiently solving LPECs using MILP reformulations. Moreover, when the solution is S-stationary, solving the MILP requires solving a single LP, complying with Proposition 3.1. This experiment suggests that solving LPECs with MILP methods as in MacMPEC can be done very efficiently in practice, both with commercial (Gurobi) and open-source (HiGHS) solvers.
Appendix D Generating a synthetic MPEC benchmark
In this section, we describe how we generate a random set of MPECs that are used in the numerical experiments of Section 5.3. The problem-generating function and code to reproduce the benchmark results is available at https://github.com/nosnoc/benchmarks/nonlinear_mpec_benchmark.
A summary of how the random MPECs are generated is given in Algorithm 4. Additionally, we describe how the problems are generated. First, we generate random LPECs as described in Hu2008 . They have the following form:
$\underset{x_{0}\in{\mathbb{R}}^{p},x_{1}\in{\mathbb{R}}^{m}}{\min}\;\quad$ $g_{0}^{\top}x_{0}+g_{1}^{\top}x_{1}$ (42a)
s.t. $Ax_{0}+Bx_{1}\geq a$ (42b)
$0\leq x_{1}\perp Mx_{0}+Nx_{1}+q\geq 0,$ (42c)
where $g_{0}\in{\mathbb{R}}^{p}$, $g_{1}\in{\mathbb{R}}^{m}$ are the cost vectors. The inequality constraint data has the following dimensions $A\in{\mathbb{R}}^{n_{\mathrm{ineq}}\times p}$, $B\in{\mathbb{R}}^{n_{\mathrm{ineq}}\times m}$, $a\in{\mathbb{R}}^{n_{\mathrm{ineq}}}$. The dimensions of the data in left hand side of (42c) are $M\in{\mathbb{R}}^{m\times p}$, $N\in{\mathbb{R}}^{m\times m}$ and $q\in{\mathbb{R}}^{m}$. A pseudocode for the generation of these LPECs can be found in (JaraMoroni2020, , p. 22). We use the same description as a starting point but allow more flexibility in changing the problem data and dimensions.
To generate a nonlinear MPEC, we make the following modification to the LPEC (42). We replace the objective in (42a) by $f(x)=f^{\mathrm{nl}}(x_{0},x_{1})+\alpha(g_{0}^{\top}x_{0}+g_{1}^{\top}x_{1})$, where $f^{\mathrm{nl}}(x_{0},x_{1})$ is a nonlinear objective test function from unconstrained problems from the CUTEst test set Gould2013 Let $c^{\mathrm{ineq}}(x)=Ax_{0}+Bx_{1}-a\geq 0$ define the inequality constraints. To introduce more nonlinearity, we consider a subset of inequality constraint indices, i.e., ${\mathcal{I}}^{\mathrm{ineq}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. Next, all inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ with $i\in{\mathcal{I}}^{\mathrm{ineq}}$ are replaced by $c_{i}^{\mathrm{ineq}}(x)+c_{i}^{\mathrm{ineq}}(x)^{2}+c_{i}^{\mathrm{ineq}}(x)^{4}\geq 0$. Moreover, we duplicated a subset of the inequality constraints (both linear and nonlinear), with indices $i\in{\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$. This makes the MPEC more degenerate as it increases the probability of violating MPEC-LICQ if inequality constraints are active. Next, we define equality constraints by introducing a slack variable $x_{2}\in{\mathbb{R}}^{m}$ for the left hand side in  (42c), which reads as $c^{\mathrm{eq}}(x)=x_{2}-Mx_{0}+Nx_{1}+q$. Analogously to the inequality constraints, we replace some constraints by $c^{\mathrm{eq}}_{i}(x)=0$ by $c^{\mathrm{eq}}_{i}(x)+c^{\mathrm{eq}}_{i}(x)^{2}+c^{\mathrm{eq}}_{i}(x)^{4}=0,i\in{\mathcal{I}}^{\mathrm{eq}}\subseteq\{1,\ldots,m\}$. Moreover, we add upper bounds $x^{\mathrm{ub}}_{0},x^{\mathrm{ub}}_{1}$ on $x_{0}$ and $x_{1}$.
The full pseudocode for generating a single problem instance is given in Algorithm 4. Thereby, $\mathcal{N}(\mu,\sigma^{2})$ denotes the normal distribution with $\mu$ mean and $\sigma$ standard deviation. The elements of $\mathcal{U}_{n}(a,b)$ are $n$ dimensional vectors whose elements are drawn from a uniform distribution on the interval $[a,b]$. Similarly, the elements of $\mathcal{U}_{n\times k}(a,b,s)$ are $n\times k$ matrices whose elements take values from a uniform distribution on $[a,b]$. The scalar $s\in\left(0,1\right]$ is the density of this random matrix, which is defined as the ratio between the number of nonzero elements and the total number of elements of this matrix. We chose the density of the random matrices to be up to 10%. For larger problems with more than 3000 variables, this results in quite dense matrices, and generating the computational graphs for the first and second-order derivatives can take several hours. To ensure that the benchmark can be solved in under a week, we set the density of a random matrix $A\in{\mathbb{R}}^{n\times m}$ to $s=\frac{n}{nm}$ for problems with more than 3000 variables.
Algorithm 4 Random MPEC generator
1:Input: Dimensions: $p,m,n_{\mathrm{ineq}}$; nonlinear objective function: $f^{\mathrm{nl}}(x_{0},x_{1})$; cost weight: $\alpha\geq 0$; proportion of nonlinear equality and inequality constraints: $s^{\mathrm{eq}},s^{\mathrm{ineq}}\in[0,1]$; proportion of duplicated inequality constraints $s^{d}\in[0,1]$.
2:Initial guess and upper bounds.
3: Generate initial guess $x_{0}^{0}\sim\mathcal{N}_{p}(0,1)$ and $x_{0}^{0}=|x_{0}^{0}|$.
4: Generate initial guess $x_{1}^{0}\sim\mathcal{N}_{m}(0,1)$ and $x_{1}^{0}=\max(0,x_{1}^{0})$.
5: Generate upper bounds $x_{0}^{\mathrm{ub}}\sim\mathcal{U}_{p}(100,1000)$ and $x_{1}^{\mathrm{ub}}\sim\mathcal{U}_{m}(100,1000)$. 6:Equality constraints.
7: Choose density for the random matrices $s\sim\mathcal{U}_{1}(0.05,0.1)$.
8: Generate random integer number $r\sim\mathrm{round}(\mathcal{U}_{1}(1,m))$.
9: Generate matrix $E\sim\mathcal{U}_{r\times(n-r)}(2,3,s)$, vectors $d_{1}\sim\mathcal{U}_{r}(0,2)$, $d_{2}\sim\mathcal{U}_{m-r}(0,2)$, and let $D_{1}=\mathrm{diag}(d_{1})$, $D_{2}=\mathrm{diag}(d_{2})$, $M=\begin{bmatrix}D_{1}&E\\ -E^{\top}&D_{2}\end{bmatrix}$.
10: Generate matrix $N\sim\mathcal{U}_{m\times p}(-1,1,s)$ and vector $q\sim\mathcal{U}_{r}(-20,-10)$,
11: Define $c^{\mathrm{eq}}(x)=Mx_{1}+Nx_{0}+q-x_{2}=0$ and $x=(x_{0},x_{1},x_{2})\in{\mathbb{R}}^{n},n=p+2m$.
12: Pick a random subset ${\mathcal{I}}^{\mathrm{eq}}\subseteq\{1,\ldots,m\}$ with cardinality $|{\mathcal{I}}^{\mathrm{eq}}|=\mathrm{round}(s^{\mathrm{eq}}m)$.
13: Replace $c_{i}^{\mathrm{eq}}(x)=0$ by $c_{i}^{\mathrm{eq}}(x)+c_{i}^{\mathrm{eq}}(x)^{2}+c_{i}^{\mathrm{eq}}(x)^{4}=0$ for all $i\in{\mathcal{I}}^{\mathrm{eq}}$. 14:Inequality constraints
15: Generate $A\sim\mathcal{U}_{n_{\mathrm{ineq}}\times p}(0,1,s)$, $B=\mathcal{U}_{n_{\mathrm{ineq}}\times m}(0,1,s)$
16: Generate $a=Ax_{0}^{0}+Bx_{1}^{0}-|\epsilon|$, $\epsilon\in\mathcal{N}_{n_{\mathrm{ineq}}}(0,1)$.
17: Define inequality constraints $c^{\mathrm{ineq}}(x)=Ax_{0}+Bx_{1}-a\geq 0$.
18: Pick a random subset ${\mathcal{I}}^{\mathrm{ineq}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$ with cardinality $|{\mathcal{I}}^{\mathrm{ineq}}|=\mathrm{round}(s^{\mathrm{ineq}}n_{\mathrm{ineq}})$.
19: Replace $c_{i}^{\mathrm{ineq}}(x)\geq 0$ by $c_{i}^{\mathrm{ineq}}(x)+c_{i}^{\mathrm{ineq}}(x)^{2}+c_{i}^{\mathrm{ineq}}(x)^{4}=0$ for all $i\in{\mathcal{I}}^{\mathrm{ineq}}$.
20: Pick a random subset ${\mathcal{I}}^{\mathrm{d}}\subseteq\{1,\ldots,n_{\mathrm{ineq}}\}$ with cardinality $|{\mathcal{I}}^{\mathrm{d}}|=\mathrm{round}(s^{\mathrm{d}}n_{\mathrm{ineq}})$ and duplicate inequality constraints $c_{i}^{\mathrm{ineq}}(x)\geq 0$ for all $i\in{\mathcal{I}}^{\mathrm{d}}$. 21:Objective function.
22: Generate $g_{0}\in\mathcal{U}_{p}(0,1)$, $g_{1}\in\mathcal{U}_{m}(1,3)$.
23: Define objective function $f(x)=f(x_{0},x_{1})+\alpha g_{0}^{\top}x_{0}+\alpha g_{1}^{\top}x_{1}$.
Our implementation allows to replacement of any random scalar variable (e.g. density $s$) by fixed values. In the implementation, the ranges of values from where the random variables are drawn can also be adapted, but for ease of notation, we write in Algorithm 4 the values used in our experiments.
To generate a test problem set, we consider 25 objective functions and create three instances of different sizes for each problem, resulting in a total of 75 problems. The the nonlinear objective functions , we use the following nonlinear test functions from Gould2013 : Fletcher, McCormick, Powell, Rosenbrock, Raydan1, Raydan2, Diagonal4, Diagonal5, Extended_Trigiaonal, Three_Exponential_Terms, Generalized_PSC1, Fletcvb3, Bdqrtic, Tridia, EG2, Edensch, Indef, Cube, Bdexp, Genhumps, Arwhead, Quartc, Cosine, NCVXQP6, DIXCHLNV. The dimensions of $p$ are unique random integers from the set $\{100,\ldots,2100\}$, $p=m$ and $n_{\mathrm{ineq}}~\mathrm{round}(\mathcal{U}_{1}(0.5p,2p))$. The remaining parameters we use in Algorithm 4 are $\alpha=10$, $s^{\mathrm{eq}}=0.01$, $s^{\mathrm{ineq}}=0.25$, and $s^{d}=0.25$.